### v2

# A priori error analysis of the proximal Galerkin method

Abstract.
The proximal Galerkin (PG) method is a finite element method for solving variational problems with inequality constraints. It has several advantages, including constraint-preserving approximations and mesh independence. This paper presents the first abstract a priori error analysis of the PG method, providing a general framework to establish convergence and error estimates. As applications of the framework, we demonstrate optimal convergence rates for both the obstacle and Signorini problems using various finite element subspaces.
Key words. Proximal Galerkin, finite element method, a priori error analysis, pointwise inequality constraint, obstacle problem, Signorini problem, Bregman proximal point.
MSC codes. 35J86, 35R35, 49J40, 65K15, 65N30.
1. Introduction
The proximal Galerkin (PG) method [36] plays a dual role, acting both as an algorithm and a discretization scheme for variational problems with pointwise inequality constraints. It is an algorithm in the sense that it comprises a sequence of operations (i.e., subproblems to be solved) leading to an approximate solution of a variational problem. It is a discretization scheme as it yields approximations that depend explicitly on chosen finite-dimensional subspaces, thereby providing a broad selection of discretization choices for the target solution.
This paper presents the first general a priori error analysis of the PG method, which has demonstrated competitive efficacy across a range of problems in applied mathematics, including classical obstacle, contact, and elastoplasticity problems [19]. Prior analyses have focused on specific discretization choices for particular problems or examined the convergence properties of the PG subproblems only after linearization [36, 28]. Instead, the present work provides a foundational advancement, dispensing with analysis of the linearized subproblems and developing a general analytical framework for PG methods applied to quadratic optimization problems in Sobolev Hilbert spaces with pointwise inequality constraints.
To illustrate the flexibility of the proposed framework, we focus on two canonical applications — the obstacle and Signorini problems — demonstrating optimal error convergence rates across a range of finite element discretizations. Key contributions of this work include the first general convergence guarantees with respect to both iteration count and mesh size, as well as the first result establishing mesh-independent iteration complexity of the PG method. These theoretical advances provide rigorous explanations for the method’s empirically observed advantages over well-established methods that preceded PG in the literature, such as the penalty and primal-dual active set methods. We refer the interested reader to [19, 40] for comparisons with these and other popular alternative methods.
1.1. Outline
The remainder of the introduction establishes the basic notation and problem setup, defines the PG method, summarizes the main results, and illustrates example problems. Section 2 defines key concepts appearing in the analytical framework, including Legendre functions and Bregman divergence. Section 3 rigorously presents the main theoretical results of the paper, including checkable conditions for existence and uniqueness of solutions to the PG subproblems, a rigorous guarantee that the objective function value will decrease monotonically with each iteration, best approximation properties for the PG solution variables, convergence rates, and asymptotic mesh-independence. Sections 4 and 5 are devoted to applications of the theory to the obstacle and Signorini problems, respectively, leading to optimal error convergence rates in each case. Finally, the paper concludes with Section 6, where we summarize our findings.
1.2. Notation
Throughout the article, we let $\Omega\subset\mathbb{R}^{n}$ ($n=1,2,3$) be an open bounded Lipschitz domain. For a given Banach space $V$, we denote by $V^{\prime}$ its topological dual space with duality pairing $\langle\cdot,\cdot\rangle$. In particular, a member $F\in V^{\prime}$ is a continuous linear functional mapping $V$ into $\mathbb{R}$, $F(v)=\langle F,v\rangle\in\mathbb{R}$. The norm $\|\cdot\|_{V^{\prime}}$ denotes the usual operator norm.
We use the standard notation for the Sobolev Hilbert spaces $H^{m}(\Omega)$ and their vector-valued counterparts $H^{m}(\Omega;\mathbb{R}^{n})$. The space $H^{1/2}(\partial\Omega)$ denotes the canonical trace space of $H^{1}(\Omega)$ functions onto the boundary $\partial\Omega$ with the quotient norm $\|\cdot\|_{H^{1/2}(\partial\Omega)}$:
$\|\hat{v}\|_{H^{1/2}(\partial\Omega)}=\inf_{\begin{subarray}{c}v\in H^{1}(\Omega)\\ \mathrm{tr}\,v=\hat{v}\end{subarray}}\|v\|_{H^{1}(\Omega)}.$
Here, $\mathrm{tr}$ denotes the trace operator. When the setting is unambiguous, we write $v|_{\partial\Omega}$ instead of $\mathrm{tr}(v)$. We also require the definition of the Lions–Magenes space on measurable $\Gamma\subset\partial\Omega$ [47]:
(1) $\widetilde{H}^{1/2}(\Gamma)=H^{1/2}_{00}(\Gamma)=\{w\in H^{1/2}(\Gamma)\mid\tilde{w}\in H^{1/2}(\partial\Omega)\},$
where $\tilde{w}$ is the extension by zero of $w$ outside of $\Gamma$; i.e., $\tilde{w}=0$ on $\partial\Omega\backslash\overline{\Gamma}$ and $\tilde{w}=w$ on $\Gamma$. We note that this space is normed by $\|w\|_{\widetilde{H}^{1/2}(\Gamma)}:=\|\tilde{w}\|_{H^{1/2}(\partial\Omega)}$ and that ${H}^{-1/2}(\Gamma)=(\widetilde{H}^{1/2}(\Gamma))^{\prime}$. For non-integer $s$, the notation $H^{s}(\Omega)$ denotes the Sobolev–Slobodeckij spaces [23, Chapter 2]. We use the notation $(\cdot,\cdot)_{\omega}$ to denote the $L^{2}(\omega)$-inner product over measurable $\omega\subset\overline{\Omega}$. If $\omega=\Omega$, we drop the subscript and denote the $L^{2}$-inner product over $\Omega$ by $(\cdot,\cdot)$.
The essential domain of a proper function $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ is given by $\operatorname{dom}f:=\{x\in\mathbb{R}^{n}:f(x)<\infty\}$. The Fréchet derivative of a mapping $F$ between normed vector spaces $X$ and $Y$ at a point $x$ is denoted by $F^{\prime}(x)$. For a linear continuous operator $B\in\mathcal{L}(U,V)$ where $U,V$ are normed vector spaces, the topological transpose (adjoint) operator $B^{\prime}\in\mathcal{L}(V^{\prime},U^{\prime})$ is defined as
(2) $\langle B^{\prime}v^{\prime},u\rangle=\langle v^{\prime},Bu\rangle\text{ for all }u\in U,v^{\prime}\in V^{\prime}.$
We consider a conforming affine shape regular simplicial mesh $\mathcal{T}_{h}$ of $\Omega$ with mesh size $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ where $h_{T}=\operatorname{diam}(T)$. Define
(3) $\mathbb{P}_{q}(\mathcal{T}_{h})=\{v\in L^{\infty}(\Omega)\mid v|_{T}\in\mathbb{P}_{q}(T)\quad~\text{ for all }T\in\mathcal{T}_{h}\},$
where $\mathbb{P}_{q}(T)$ denotes the space of polynomials of total degree less than or equal to $q$ on $T$. Denote by $\mathcal{N}_{h}$ the set of element vertices (nodes) in $\mathcal{T}_{h}$ and by $\{\varphi_{z}\}_{z\in\mathcal{N}_{h}}$ the associated nodal basis functions of polynomial degree 1. The collection of $n+1$ vertices of an element $T\in\mathcal{T}_{h}$ is denoted by $\mathcal{V}_{T}$. For each $z\in\mathcal{N}_{h}$, let $\omega_{z}$ denote the union of elements sharing the node $z$.
For constants $a$ and $b$, we use the standard notation $a\lesssim b$ whenever there exists a constant $c$ that depends on neither the mesh size $h$ nor on the proximal point parameters such that $a\leq cb$.
1.3. General setup
The PG method is versatile and has been successfully applied to a diverse set of variational problems with inequality constraints [19]. In this work, we consider constrained optimization problems of the following form
(4) $\min_{v\in K}E(v),$
where $E\colon V\rightarrow\mathbb{R}$ is an energy function and $K$ is a closed, convex, and non-empty set taking the general form
(5) $K=\{v\in V\mid Bv(x)\in C(x)\text{ for almost every }x\in\Omega_{d}\subset\overline{\Omega}\}.$
Here, $V$ is a given (affine) Hilbert space, $\Omega_{d}$ is a Hausdorff-measurable set with Hausdorff dimension $d\leq n$ and Hausdorff measure $\operatorname{d}\!\mathcal{H}_{d}$, $B\colon V\rightarrow Q$ is a surjective bounded linear map, whose image $Q=\operatorname{im}B$ is continuously and densely embedded in $L^{2}(\Omega_{d};\mathbb{R}^{m})$, and $C(x)\subset\mathbb{R}^{m}$, which may vary with $x$, is a closed convex set with a nonempty interior.
In the definition of the feasible set $K$, the map $B$ can be understood to define observables $o\in Q$ that are restricted pointwise a.e. to $C(x)$. For example, in obstacle problems, where $V\subset H^{1}(\Omega)$ and functions $v\in K$ satisfy $v\geq\phi$ a.e. in $\Omega$, we have $\Omega_{d}=\Omega$, $B=\mathrm{id}$ (the identity operator), and $C(x)=[\phi(x),\infty)$ for a given obstacle $\phi\in L^{\infty}(\Omega)$. In this case, the observables are simply the unknown functions $v\in V$. For contact problems, $\Omega_{d}=\Gamma$ is a Hausdorff-measurable subset of $\partial\Omega$, and the observables are the normal traces of the functions $v\in V$ restricted to $\Gamma$; namely, $o=v|_{\Gamma}\cdot n$. Refer to [19] as well as Examples 1.1 and 1.2, given below, for more details. For notational convenience later on, we introduce specific notation for the set of constrained observables:
(6) $\mathcal{O}=\{o\in L^{2}(\Omega_{d};\mathbb{R}^{m})\mid o(x)\in C(x)\text{ for almost every }x\in\Omega_{d}\subset\overline{\Omega}\}.$
Throughout this work, we consider only quadratic energies,
(7) $E(u)=\frac{1}{2}a(u,u)-F(u),$
where $a$ is a symmetric, continuous, and coercive bilinear form over the space $V$, satisfying
(8) $a(u,u)\geq\nu\|u\|_{V}^{2},\quad a(u,v)\leq M\|u\|_{V}\|v\|_{V}~\text{ for all }u,v\in V,$
for positive constants $\nu$ and $M$, and $F$ is a bounded linear functional on $V$. The convexity of $K$ implies that the model problem 4 is equivalent to the following variational inequality [16, Theorem 6.1-2]: find $u^{*}\in K$, such that
(9) $a(u^{*},v-u^{*})\geq F(v-u^{*})~\text{ for all }v\in K.$
Owing to the coercivity of $a$, 9 admits a unique solution [10, Theorem 5.6]. Moreover, $E(u^{\ast})\leq E(v)$ for all $v\in K$.
1.4. The proximal Galerkin method
We present the method here and refer to Section 3 for more details on its derivation. Consider two conforming discrete subspaces $V_{h}\subset V$ and $W_{h}\subset W:=L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. The PG method, given in Algorithm 1, consists of iteratively solving for primal solutions $u_{h}^{k}\in V_{h}$ and latent solutions $\psi_{h}^{k}\in W_{h}$. Note that some form of Newton’s method is usually used to solve each subproblem in practice; see, e.g., the implementations in [20].
Algorithm 1 The Proximal Galerkin Method
1:input: Initial latent solution guess $\psi_{h}^{0}\in W_{h}$, a sequence of positive proximity parameters $\{\alpha_{k}\}$, and a functional $\mathcal{R}^{*}$ with $\nabla\mathcal{R}^{*}:W\rightarrow\mathcal{O}$.
2:Initialize $k=1$.
3:repeat
4: Find $u_{h}^{k}\in V_{h}$ and $\psi_{h}^{k}\in W_{h}$ such that
(10a) $\alpha_{k}\,a(u^{k}_{h},v_{h})+b(v_{h},\psi^{k}_{h}-\psi_{h}^{k-1})$ $=\alpha_{k}\,F(v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(10b) $b(u^{k}_{h},w_{h})-(\nabla\mathcal{R}^{*}(\psi^{k}_{h}),w_{h})_{\Omega_{d}}$ $=0$  $~\text{ for all }w_{h}\in W_{h}.$
5: Assign $k\leftarrow k+1$.
6:until a convergence test is satisfied.
In Algorithm 1, $b:V\times W\rightarrow\mathbb{R}$ is a bilinear form corresponding to the operator $B$ in the feasible set 5. Namely,
(11) $b(v,w)=(Bv,w)_{\Omega_{d}}~\text{ for all }v\in V,\,w\in W.$
The map $\nabla\mathcal{R}^{*}$ is the inverse of the Fréchet derivative of a suitably chosen Legendre function $\mathcal{R}$; see Section 2.1 and Example 2.1 for more details. We refer to Algorithms 2 and 3, given in the sections below, for applications of this algorithm to the obstacle and Signorini problems, respectively, with particular choices of $\mathcal{R}$.
The saddle point system 10 also produces a non-polynomial approximation of the observable $o^{\ast}=Bu^{\ast}$:
(12) $o_{h}^{k}=\nabla\mathcal{R}^{*}(\psi_{h}^{k}),\quad k\geq 0.$
This variable is always constraint-preserving because $\operatorname{im}\nabla\mathcal{R}^{*}\subset\mathcal{O}$. Likewise, $o_{h}^{k}(x)\in C(x)$ for all $x\in\Omega_{d}$. In addition, we define the dual variables
(13) $\lambda^{k}_{h}=(\psi_{h}^{k-1}-\psi_{h}^{k})/\alpha_{k},\quad k\geq 1,$
which are viewed as $\lambda_{h}^{k}\in Q^{\prime}$ via $\langle\lambda_{h}^{k},q\rangle=(\lambda_{h}^{k},q)_{\Omega_{d}}$. As we show below, these dual variables converge to the unique $\lambda^{*}\in Q^{\prime}$ satisfying
(14) $B^{\prime}\lambda^{*}=E^{\prime}(u^{*})\text{ in }V^{\prime}.$
1.5. Main results
  • We prove that every PG subproblem is well-defined for the general setup of Sections 1.3 and 1.4. More precisely, Theorem 3.1 establishes the existence and uniqueness of solutions to the discrete nonlinear subproblems 10 provided certain compatibility conditions between the subspaces $V_{h}$ and $W_{h}$ are satisfied. Additionally, we prove important new energy dissipation and stability estimates for 10.
  • We provide a general framework for the error analysis of Algorithm 1. This framework shows that the existence and optimality of certain enriching and Fortin operators, defined below, are sufficient to derive error estimates and mesh-independence results; see Theorem 3.3, Theorem 3.6, and Corollary 1.
  • We demonstrate applications of this framework to the analysis of obstacle and Signorini problems; see Section 4 and Section 5. In particular, we construct and prove error rates for the Fortin and enriching operators used in the proposed framework of Section 3.
  • Finally, one of the main results of this paper may be summarized by the following estimate. If $V\subset H^{1}(\Omega)$ and the solution $u^{*}\in H^{1+s}(\Omega)$ with $E^{\prime}(u^{*})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{r-1}(\Omega)}$ for some $s,r\in(0,1]$, then
    (15) $\|u^{*}-u_{h}^{\ell}\|^{2}_{H^{1}(\Omega)}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{Q^{\prime}}\leq\frac{C_{\rm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}h^{2\cdot\min\{r,s\}}$
    for all $\ell\geq 1$ and $h>0$, where the constants $C_{\rm{stab}}$ and $C_{\rm{reg}}$ are independent of $h$, $\{\alpha_{k}\}$, and $\ell$. We first state 15 in Theorem 3.6, where we prove it under general assumptions. We then prove 15 again in Corollaries 2 and 3 for the obstacle and Signorini problems, respectively, by verifying the general assumptions for specific choices of $V_{h}$ and $W_{h}$.
1.6. Example problems
We provide three examples illustrating the setup. The forthcoming sections expand on the first two examples. The PG method has not yet been applied to the third example, and so further analysis is reserved for follow-up work. Together, these examples illustrate the need for a theory that comprises arbitrary Hilbert spaces $V$, observation maps $B$, subsets $\Omega_{d}$, and convex sets $C$. Numerical experiments and additional examples can be found in [19].
Example 1.1 (Obstacle problem).
For the obstacle problem, set the space $V=H^{1}_{0}(\Omega)$. One seeks to minimize the Dirichlet energy
(16) $E(u)=\frac{1}{2}\|\nabla u\|^{2}_{L^{2}(\Omega)}-\int_{\Omega}fu\operatorname{d}\!x,$
over the feasible set
(17) $K=\{v\in H^{1}_{0}(\Omega)\mid v\geq\phi\text{ a.e.\ in }\Omega\},$
where $f\in L^{2}(\Omega)$ and $\phi\in L^{\infty}(\Omega)$ is a given obstacle satisfying $\phi\leq 0$ a.e. on $\partial\Omega$. The forms $a:H^{1}(\Omega)\times H^{1}(\Omega)\rightarrow\mathbb{R}$ and $F\colon H^{1}(\Omega)\rightarrow\mathbb{R}$ are given by
(18) $a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v\operatorname{d}\!x,\quad F(v)=\int_{\Omega}fv\operatorname{d}\!x.$
Note that $a$ is coercive over $V$ due to Poincaré’s inequality. To view $K$ in the general form 5, set $B$ to be the identity operator on $V=Q$, take $C(x)=[\phi(x),\infty)$, and let $\Omega_{d}=\Omega$. In this case, $\lambda^{*}=E^{\prime}(u^{*})\in V^{\prime}$ with $\langle E^{\prime}(u^{*}),\cdot\rangle=a(u^{*},\cdot)-F(\cdot)$.
Example 1.2 (Signorini problem).
We consider disjoint boundaries $\Gamma_{\mathrm{T}}$ and $\Gamma_{\mathrm{D}}$ which are measurable subsets of $\partial\Omega$ with $\partial\Omega=\overline{\Gamma_{\mathrm{D}}\cup\Gamma_{\mathrm{T}}}$. The Signorini problem consists of finding the displacement $u\in V=(H^{1}_{\mathrm{D}}(\Omega))^{n},$ where $H^{1}_{\mathrm{D}}(\Omega):=\{v\in H^{1}(\Omega)\mid v|_{\Gamma_{\mathrm{D}}}=0\}$ minimizing the strain energy function
(19) $E(u)=\frac{1}{2}\int_{\Omega}\mathsf{C}\,\epsilon(u):\epsilon(u)\operatorname{d}\!x-\int_{\Omega}f\cdot u\operatorname{d}\!x,$
over the convex and closed set
(20) $K=\{u\in V\mid u\cdot n\leq g\text{ a.e.\ on }\Gamma_{\mathrm{T}}\}.$
Here, $\epsilon(u)=\frac{1}{2}(\nabla u+\nabla u^{\top})$ is the linearized strain tensor, $\mathsf{C}\colon\mathbb{R}^{n\times n}_{\mathrm{sym}}\rightarrow\mathbb{R}^{n\times n}_{\mathrm{sym}}$ is a symmetric positive definite material tensor, $f\in L^{2}(\Omega;\mathbb{R}^{n})$, $n$ is the unit outward normal vector on $\partial\Omega$, and $g\in L^{\infty}(\Gamma_{\mathrm{T}})$ with $g\geq 0$ is a prescribed gap function.
We can write the set $K$ in the form 5 by setting $Bu=-u|_{\Gamma_{\mathrm{T}}}\cdot n$, $Q=\widetilde{H}^{1/2}(\Gamma_{\mathrm{T}})$, $\Omega_{d}=\Gamma_{\mathrm{T}}$, and $C(x)=[-g(x),\infty)$. The bilinear form $a:H^{1}(\Omega;\mathbb{R}^{n})\times H^{1}(\Omega;\mathbb{R}^{n})\rightarrow\mathbb{R}$ and linear form $F\colon H^{1}(\Omega;\mathbb{R}^{n})\rightarrow\mathbb{R}$ are given by
(21) $a(u,v)=\int_{\Omega}\mathsf{C}\,\epsilon(u):\epsilon(v)\operatorname{d}\!x,\quad F(v)=\int_{\Omega}f\cdot v\operatorname{d}\!x.$
Korn’s inequality guarantees the coercivity of the bilinear form $a$ over $V$; see, e.g., [25, Theorem 42.9]. In this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n\in{H}^{-1/2}(\Gamma_{\mathrm{T}})=Q^{\prime}$.
Example 1.3 (Image restoration).
Fix $f\in L^{2}(\Omega)$ and $\beta>0$. It is well-known [37] that the pre-dual of the classical total bounded variation-regularized tracking problem [43],
$\text{minimize }~~J(u):=|Du|(\Omega)+\frac{\beta}{2}\int_{\Omega}(u-f)^{2}\operatorname{d}\!x~\text{ over }u\in BV(\Omega)\cap L^{2}(\Omega)\,$,
where $BV(\Omega)$ is the space of functions of bounded variation over $\Omega$ [4, 26] and
$|Du|(\Omega)=\sup\bigg\{\int_{\Omega}u\operatorname{div}\phi\operatorname{d}\!x\mid\phi\in C^{1}_{c}(\Omega;\mathbb{R}^{n}),~|\phi|_{\ell^{\infty}}\leq 1\text{ a.e.\ in }\Omega\bigg\}$
denotes the $BV(\Omega)$-seminorm, can be expressed as a bilaterally constrained optimization problem. More specifically, we are interested in finding a unique $u^{\ast}\in BV(\Omega)\cap L^{2}(\Omega)$ such that $J(u^{\ast})\leq J(v)$ for all $v\in BV(\Omega)\cap L^{2}(\Omega)$. This problem is well-posed, and in [37] it is shown that its solution satisfies the following identity:
$u^{\ast}=f+\beta^{-1}\operatorname{div}p^{\ast},$
where $p^{\ast}\in H_{0}(\operatorname{div},\Omega)=\{p\in L^{2}(\Omega;\mathbb{R}^{n})\mid\operatorname{div}p\in L^{2}(\Omega),~p|_{\partial\Omega}\cdot n=0\}$ is the unique minimizer of the energy function
(22) $E(p)=\frac{1}{2}\int_{\Omega}(\operatorname{div}p)^{2}\operatorname{d}\!x+\frac{\gamma}{2}\int_{\Omega}(\operatorname{proj}p)^{2}\operatorname{d}\!x+\beta\int_{\Omega}f\operatorname{div}p\operatorname{d}\!x$
over the convex set
$K=\big\{p\in H_{0}(\operatorname{div},\Omega)\mid|p_{i}|\leq 1\text{ a.e.\ in }\Omega\text{ for each }i=1,\ldots,n\big\}.$
In 22, $\gamma>0$ is a fixed parameter and $\operatorname{proj}$ is the orthogonal projection $H_{0}(\operatorname{div},\Omega)\to\{p\in H_{0}(\operatorname{div},\Omega)\mid\operatorname{div}p=0\}$.
We can write the set $K$ in the form of 5 by taking $B$ to be the identity on $V=Q=H_{0}(\operatorname{div},\Omega)$, $\Omega_{d}=\Omega$, and $C=[-1,1]^{n}$. In this case, the bilinear form $a\colon V\times V\rightarrow\mathbb{R}$ and linear form $F\colon V\rightarrow\mathbb{R}$ are given by
(23) $a(p,q)=\int_{\Omega}\operatorname{div}p\,\operatorname{div}q\,\operatorname{d}\!x+\gamma\int_{\Omega}\operatorname{proj}p\,\operatorname{proj}q\,\operatorname{d}\!x,\quad F(q)=-\beta\int_{\Omega}f\operatorname{div}q\,\operatorname{d}\!x.$
Note that $a$ is coercive over the Hilbert space $V$ due to a Friedrichs’ inequality; see, e.g., [18, Lemma 2.8].
1.7. Closed observation maps: A conjecture
Some important problems with feasible sets of the form 5 do not fit into the general setup described in Section 1.3 because the observation map $B$ does not map onto a dense subset of $L^{2}(\Omega_{d};\mathbb{R}^{m})$. For instance, consider the classical elastoplastic torsion problem [48, 9], which involves minimizing the Dirichlet energy 16 over a feasible set with gradient constraints, such as
(24) $K=\{v\in H^{1}_{0}(\Omega)\mid|\nabla v|\leq 1\text{ a.e.\ in }\Omega\}.$
Note that this set is recovered from 5 by setting $C$ to be the closed unit ball in $\mathbb{R}^{n}$, $V=H^{1}_{0}(\Omega)$, $B=\nabla$, and $\Omega_{d}=\Omega$. For further details, see [19, Example 6]. If $n\geq 2$, then $\operatorname{im}B$ is a closed proper subspace of $L^{2}(\Omega;\mathbb{R}^{n})$.
Informed by numerical experiments in [19, 40], we conjecture that the PG iterates $u_{h}^{k}$ can also be shown to converge to the exact solution of 4 if $B\colon V\to L^{2}(\Omega_{d};\mathbb{R}^{m})$ is a closed operator. However, unlike the analysis below, we can not treat each subproblem 10 as a singularly-perturbed nonlinear saddle-point problem in such a setting, and we must also account for the possibly non-trivial kernel of $B^{\prime}$. In turn, we expect different general results with such observation maps, and we do not consider them further in this work.
2. Preliminaries
We briefly recall two key concepts fundamental to the derivation and analysis of PG methods: Legendre functions and Bregman divergences. We refer the interested reader to [41, 11, 6, 36] for more details. We then derive the PG method.
2.1. Legendre functions
Algorithm 1 depends on the specific choice of the functional $\mathcal{R}^{*}$, which satisfies $\nabla\mathcal{R}^{*}:L^{\infty}(\Omega_{d},\mathbb{R}^{m})\rightarrow\mathcal{O}$. The construction of $\mathcal{R}^{*}$ relies on the concept of a Legendre function [42]. In this work, a function $L:\mathbb{R}^{m}\rightarrow\mathbb{R}\cup\{+\infty\}$ is called a Legendre function if it is proper with $\operatorname{int}(\operatorname{dom}L)\neq\emptyset$, strictly convex and differentiable on $\operatorname{int}(\operatorname{dom}L)$ with a singular gradient on the boundary of $\operatorname{dom}L$. This subsection aims to briefly show how Legendre functions are utilized to define $\mathcal{R}^{*}$ as the convex conjugate of a superposition operator $\mathcal{R}$.
Consider a Carathéodory function $R:\Omega_{d}\times\mathbb{R}^{m}\rightarrow\mathbb{R}\cup\{+\infty\}$ where $R(x,\cdot)$ is a Legendre function with $\operatorname{dom}(R(x,\cdot))=C(x)$ f.a.e. $x\in\Omega_{d}$. Let
$\mathcal{R}(w)(x)=R(x,w(x)),\quad x\in\Omega_{d},\,w\in L^{2}(\Omega_{d}),$
be the corresponding superposition operator with
$\nabla\mathcal{R}(u)(x)=\partial_{u}R(x,u(x)).$
Assumption 1 (Continuity).
We assume that $\mathcal{R}:\mathcal{O}\rightarrow L^{1}(\Omega_{d})$ is continuous; i.e., if $\{o_{n}\},\;o\in\mathcal{O}$, and $\lim_{n\rightarrow\infty}\|o_{n}-o\|_{L^{2}(\Omega_{d})}\rightarrow 0$, then $\lim_{n\rightarrow\infty}\|\mathcal{R}(o_{n})-\mathcal{R}(o)\|_{L^{1}(\Omega_{d})}=0$.
The convex conjugate of $R(x,\cdot)$ and its associated superposition operator are given by
(25) $R^{*}(x,z)=\sup_{y\in\mathbb{R}}\big\{zy-R(x,y)\big\},\,\;\;\mathcal{R}^{*}(\psi)(x)=R^{*}(x,\psi(x))\,.$
In the sequel, we tacitly assume a supercoercivity property of the map $R(x,\cdot)$; namely, $R(x,y)/|y|\rightarrow\infty$ as $|y|\rightarrow\infty$ for a.e. $x\in\Omega$. This ensures that $R^{*}(x,\cdot)$ is well-defined and continuously differentiable over all of $\mathbb{R}^{m}$ [6, Proposition 2.16]; see also [42, Corollary 13.3.1]. Likewise, we can conclude that $\nabla\mathcal{R}^{*}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. We now utilize the following relation, first demonstrated for Legendre functions in [41]:
(26) $\nabla\mathcal{R}^{*}=(\nabla\mathcal{R})^{-1}.$
In turn, we conclude that $\operatorname{dom}(\nabla\mathcal{R})=\operatorname{im}(\nabla\mathcal{R}^{*})\subset\mathcal{O}$, implying $\nabla\mathcal{R}^{*}(\psi)(x)\in\operatorname{int}C(x)$ f.a.e. $x\in\Omega_{d}$. We provide three explicit examples in Examples 2.1, 2.2 and 2.3 below, and we refer to [19, Table 1] for more.
Identity 26 also allows us to define a latent representation of every observable $Bu\in\operatorname{dom}(\nabla\mathcal{R})$; namely,
(27) $\psi=\nabla\mathcal{R}(Bu)\iff\nabla\mathcal{R}^{*}(\psi)=Bu.$
We refer to such functions $\psi\colon\Omega_{d}\to\mathbb{R}^{m}$ as latent variables.
In the subsequent sections, we will make use of the following identity, which can be derived by directly expressing the convex conjugate $R^{*}(x,z)$ given in 25 by
(28) $R^{*}(x,z)=z\left(\partial_{u}R(x,\cdot)\right)^{-1}(z)-R(x,\left(\partial_{u}R(x,\cdot)\right)^{-1}(z)).$
Likewise, we deduce that
(29) $\mathcal{R}^{*}(\psi)=\psi Bu-\mathcal{R}(Bu)$
for any $u$ and $\psi$ satisfying 27.
Example 2.1 (Shannon entropy).
Consider Example 1.1 and define
$R(x,y)=(y-\phi(x))\ln(y-\phi(x))-(y-\phi(x)),$
if $y\geq\phi(x)$ and $R(x,y)=+\infty$ otherwise. The corresponding superposition operator is
$\mathcal{R}(u)=(u-\phi)\ln(u-\phi)-(u-\phi).$
We deduce that $\nabla\mathcal{R}(u)=\ln(u-\phi)$ whenever $u\in\operatorname{dom}(\nabla\mathcal{R})$ where
$\operatorname{dom}(\nabla\mathcal{R})=\{u\in L^{\infty}(\Omega)\mid\operatorname{ess}\operatorname{inf}(u-\phi)>0\}.$
The continuity of $\mathcal{R}$ on $\mathcal{O}$ (Assumption 1) follows from [36, Theorem 4.1]. A simple computation shows that $R^{*}(x,z)=\exp(z)+\phi(x)z$ with
(30) $\mathcal{R}^{*}(\psi)=\exp(\psi)+\phi\psi,\quad\quad\nabla\mathcal{R}^{*}(\psi)=\exp(\psi)+\phi.$
Note that $\nabla\mathcal{R}^{*}$ is well defined for all of $L^{\infty}(\Omega)$ and $\nabla\mathcal{R}^{*}(\psi)>\phi$ a.e. in $\Omega$.
Example 2.2 (Fermi–Dirac binary entropy).
For bilateral constraints, $\underline{u}(x)\leq Bu(x)\leq\overline{u}(x)$ in $\Omega_{d}$, we define
(31) $R(x,y)=(y-\underline{u}(x))\ln(y-\underline{u}(x))+(\overline{u}(x)-y)\ln(\overline{u}(x)-y),$
if $\underline{u}(x)\leq y\leq\overline{u}(x)$ and $R(x,y)=+\infty$ otherwise. The corresponding superposition operator $\mathcal{R}$ is continuous on $\mathcal{O}$; i.e., it satisfies Assumption 1 [36, Lemma 3.2]. Here, $\nabla\mathcal{R}(o)=\ln(o-\underline{u})-\ln(\overline{u}-o)$ with
$\operatorname{dom}(\nabla\mathcal{R})=\{o\in L^{\infty}(\Omega_{d})\mid\operatorname{ess}\operatorname{inf}(o-\underline{u})>0\text{ and }\operatorname{ess}\operatorname{sup}(\overline{u}-o)<0\}.$
With 26, we derive that
(32) $\nabla\mathcal{R}^{*}(\psi)=\frac{\underline{u}+\overline{u}\exp(\psi)}{1+\exp(\psi)}.$
Observe that $\underline{u}<\nabla\mathcal{R}^{*}(\psi)<\overline{u}$. This entropy functional is suitable for Example 1.3 with $\underline{u}=-1$ and $\overline{u}=1$ where the latent variable space is vector-valued, $W=L^{\infty}(\Omega;\mathbb{R}^{n})$.
Example 2.3 (Hellinger entropy).
For Euclidean norm constrains $|Bu|\leq\gamma$, cf. 24, one can select the Hellinger entropy
$\mathcal{R}(u)=-\sqrt{\gamma^{2}-|u|^{2}},\text{ with }\nabla\mathcal{R}^{*}(\psi)=\frac{\gamma}{\sqrt{1+|\psi|^{2}}}\psi.$
In this case, we have that
$\nabla\mathcal{R}(o)=\frac{o}{\sqrt{\gamma^{2}-|o|^{2}}},\quad\operatorname{dom}(\nabla\mathcal{R})=\{o\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})\mid\operatorname{ess}\operatorname{sup}(|o|-\gamma)<0\}.$
This entropy has been used in the PG framework for gradient norm constraints, see Section 1.7 and [19, Example 6].
2.2. Bregman divergences
The Legendre functions defined in Section 2.1 allow one to define the Bregman divergence, which is a key ingredient in the derivation of generalized proximal point methods [11, 36]. For $u\in\rm{dom}(\mathcal{R})$ and $v\in\mathrm{dom}(\nabla\mathcal{R})$, the Bregman divergence is given by the error in the first order Taylor expansion of an associated convex function $\mathcal{R}$:
(33) $\mathcal{D}(u,v)=\mathcal{R}(u)-\mathcal{R}(v)-\nabla\mathcal{R}(v)(u-v).$
Observe that $\mathcal{D}(u,v)\geq 0$ and $\mathcal{D}(v,v)=0$. We will also use the dual or conjugate divergence
(34) $\mathcal{D}^{*}(\eta,\psi)=\mathcal{R}^{*}(\eta)-\mathcal{R}^{*}(\psi)-\nabla\mathcal{R}^{*}(\psi)(\eta-\psi).$
The Bregman divergence and its conjugate are linked as follows. If $\eta=\nabla\mathcal{R}(v)$ and $\psi=\nabla\mathcal{R}(u)$ for $u,v\in\mathrm{dom}(\nabla\mathcal{R})$, then
(35) $\mathcal{D}^{*}(\eta,\psi)=\mathcal{D}(u,v).$
The proof can be found in [2], see also [6, Theorem 3.9]. We also recall the three points identity [13, Lemma 3.1]:
(36) $\mathcal{D}(u,v)-\mathcal{D}(u,w)+\mathcal{D}(v,w)$ $=(\nabla\mathcal{R}(v)-\nabla\mathcal{R}(w))(v-u).$
The same identity holds for $\mathcal{D}^{*}$ with $\mathcal{R}^{*}$ replacing $\mathcal{R}$.
2.3. Deriving the proximal Galerkin method
The PG method can be seen as a conforming finite element discretization of a continuous-level algorithm known as the latent variable proximal point algorithm (LVPP) [19]. However, LVPP is itself just a convenient rewriting of the Bregman proximal point algorithm [13]:
(37) $u^{k}=\operatorname*{arg\,min}_{u\in K}E(u)+\alpha_{k}^{-1}\int_{\Omega_{d}}\mathcal{D}(Bu,Bu^{k-1})\operatorname{d}\!\mathcal{H}_{d},\qquad k=1,2,\ldots$
The algorithm 37 leverages the Legendre function $\mathcal{R}$ to adaptively regularize 4, resulting in iterates $u^{k}$ that converge at a controllable speed to the global minimizer $u^{\ast}$. The message behind the LVPP reformulation is that the subproblems in 37 are easy to discretize and solve if the latent variable $\psi$ in 27 is incorporated.
Formally, choosing $R(x,\cdot)$ with singular derivatives at $\partial C(x)$ f.a.e. $x\in\Omega_{d}$ (cf. Section 2.1), one expects that $Bu^{k}\in\operatorname{im}\nabla\mathcal{R}^{*}$; in particular, $Bu^{k}(x)\in\mathrm{int}\,C(x)$ f.a.e. $x\in\Omega_{d}$. Likewise, the solutions $u^{k}$ of the regularized subproblems 37 are characterized by variational equations:
(38) $\text{find }u^{k}\in K\text{ such that }~\alpha_{k}\langle E^{\prime}(u^{k}),v\rangle+(\nabla\mathcal{R}(Bu^{k}),Bv)_{\Omega_{d}}=(\nabla\mathcal{R}(Bu^{k-1}),Bv)_{\Omega_{d}}$
for all $v\in V$. Introducing the latent variables $\psi^{k}=\nabla\mathcal{R}(Bu^{k})$ (i.e., $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})$ by 26) to rewrite the resulting equations in saddle-point form yields the LVPP algorithm, and discretizing the resulting saddle-point problems leads to the PG method [19]. We refer the reader to [36] for a detailed derivation of LVPP for the obstacle problem and to [28, Section 3] for a brief summary.
The LVPP algorithm reads as follows: for some starting point $\psi^{0}\in{W}$ and an unsummable sequence of positive parameters $\{\alpha_{k}\}$, find ($u^{k},\psi^{k})\in V\times W$ such that
(39a) $\alpha_{k}a(u^{k},v)+b(v,\psi^{k}-\psi^{k-1})$ $=\alpha_{k}F(v)$  $~\text{ for all }v\in V,$
(39b) $b(u^{k},w)-(\nabla\mathcal{R}^{*}(\psi^{k}),w)_{\Omega_{d}}$ $=0$  $~\text{ for all }w\in W,$
and $k=1,\ldots$ Note that from 39b, the latent variables $\psi^{k}$ satisfy a crucial identity: $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})\in\mathcal{O}$. Proving that 39 is well-posed is generally a challenging task. To date, it has only been accomplished for the obstacle problem (cf. Example 1.1) under suitable regularity assumptions on the problem data [36]. Fortunately, in the a priori error analysis that follows, we do not rely on the continuous-level subproblems 39 in any way. Instead, we focus solely on the discretized subproblems 10.
3. Framework for the analysis of the proximal Galerkin method
This section presents our main results in a general framework. We first introduce a compatible subspace condition to demonstrate that the PG method is well-defined. We then show that PG is endowed with a convenient energy decay property, leading to best approximation results and error convergence rates if a so-called enriching operator exists. The assumptions of this section are verified for the obstacle and Signorini problems in Section 4 and Section 5, respectively.
3.1. Compatible subspaces
A critical condition for our analysis is that the finite-dimensional subspaces $V_{h}\subset V$ and $W_{h}\subset W\subset Q^{\prime}$ satisfy the discrete inf-sup or Ladyzhenskaya–Babuška–Brezzi (LBB) condition
(40) $\inf_{w\in W_{h}}\sup_{v\in V_{h}}\frac{b(v,w)}{\|v\|_{V}\|w\|_{Q^{\prime}}}=\beta_{h}>\beta_{0},$
where $\beta_{0}>0$ is a mesh-independent positive constant. This condition is closely related to the continuous inf-sup condition
(41) $\inf_{w\in W}\sup_{v\in V}\frac{b(v,w)}{\|v\|_{V}\|w\|_{Q^{\prime}}}=\beta>0.$
The density of $W$ in $Q^{\prime}$ and the closed range theorem can be used to show that 41 implies $B^{\prime}\colon Q^{\prime}\to V^{\prime}$ is bounded from below. Likewise, 40 ensures that $B^{\prime}$ remains bounded from below (uniformly in $h$) after discretization.
To prove 40, it suffices to exhibit a continuous so-called Fortin operator $\Pi_{h}\colon V\rightarrow V_{h}$ satisfying $\|\Pi_{h}v\|_{V}\lesssim\|v\|_{V}$ for all $v\in V$ and
(42) $b(v-\Pi_{h}v,w_{h})=0~\text{ for all }v\in V,\;w_{h}\in W_{h};$
see, e.g., [25, Lemma 26.9] and [7, Section 5.4.3]. In what follows, we let $\|\Pi_{h}\|$ denote the operator norm of $\Pi_{h}$.
We consider the following examples to further illustrate the inf-sup conditions.
Example 3.1 (The obstacle problem, part 2).
Recall that $V=H^{1}_{0}(\Omega)$ with norm $\|\cdot\|_{V}=|\cdot|_{H^{1}(\Omega)}$ and $Q=H^{1}_{0}(\Omega)$. The norm $\|\cdot\|_{Q^{\prime}}$ is the $H^{-1}(\Omega)$ norm. For $w\in W=L^{\infty}(\Omega)$, we have
(43) $\|w\|_{H^{-1}(\Omega)}=\sup_{v\in H^{1}_{0}(\Omega)}\frac{|\int_{\Omega}wv\operatorname{d}\!x|}{\|\nabla v\|_{L^{2}(\Omega)}},$
and the bilinear form $b$ is simply the $L^{2}(\Omega_{d})$-inner product, given by
(44) $b(v,w)=(v,w)_{\Omega_{d}}=\int_{\Omega}vw\operatorname{d}\!x~\text{ for all }v,w\in L^{2}(\Omega).$
The associated LBB condition 41 holds with equality for $\beta=1$, which immediately follows from the definition of the $H^{-1}(\Omega)$ norm. In Section 4, we provide two examples of compatible subspaces $V_{h}\times W_{h}$ satisfying 40 with their corresponding Fortin operators.
Example 3.2 (The Signorini problem, part 2).
Recall that $V=(H_{\mathrm{D}}^{1}(\Omega))^{n}$. The bilinear form $b$ is given by
(45) $b(v,w)=\int_{\Gamma_{\mathrm{T}}}v\cdot n\,w\operatorname{d}\!s~\text{ for all }v\in(H_{\mathrm{D}}^{1}(\Omega))^{n},~w\in L^{2}(\Gamma_{\mathrm{T}}).$
The norm $\|\cdot\|_{Q^{\prime}}$ is defined as follows:
(46) $\|w\|_{Q^{\prime}}=\|w\|_{H^{-1/2}(\Gamma_{\mathrm{T}})}=\sup_{\hat{v}\in\widetilde{H}^{1/2}(\Gamma_{\rm T})}\frac{|\int_{\Gamma_{\mathrm{T}}}w\hat{v}\operatorname{d}\!s|}{\|\hat{v}\|_{\widetilde{H}^{1/2}(\Gamma_{\rm T})}}$
The following inf-sup condition holds with a constant $\beta>0$ [14, Proposition 7.2 and Remark 7.2]:
(47) $\inf_{w\in L^{\infty}(\Gamma_{\mathrm{T}})}\sup_{v\in(H^{1}(\Omega_{\mathrm{D}}))^{n}}\frac{\int_{\Gamma_{\mathrm{T}}}v\cdot nw\operatorname{d}\!s}{\|v\|_{H^{1}(\Omega)}\|w\|_{H^{-1/2}(\Gamma_{\mathrm{T}})}}\geq\beta.$
In Section 5, we provide an example of $V_{h}\times W_{h}$ satisfying 40, and we construct the corresponding Fortin operator.
Example 3.3 (Point-wise divergence constraints).
The present framework also allows one to handle a limited number of cases where $B$ is a differential operator; cf. Section 1.7. For example, consider
$K=\{v\in H(\mathrm{div},\Omega)\mid|\nabla\cdot v|\leq 1\text{ a.e.\ in }\Omega\},$
and define the energy functional $E(v)=\frac{1}{2}\|v\|^{2}_{H(\mathrm{div},\Omega)}-(f,v)$ for all $v\in V=H(\mathrm{div},\Omega)$ and some fixed $f\in L^{2}(\Omega;\mathbb{R}^{n})$. The form $b:V\times Q\rightarrow\mathbb{R}$ then reads
$b(v,w)=\int_{\Omega}\nabla\cdot v\,w\operatorname{d}\!x,$
with $Q=L^{2}(\Omega)$. For the Legendre function $\mathcal{R}$, one can choose the Fermi–Dirac entropy given in Example 2.2 with $\underline{u}=-1$ and $\overline{u}=1$, although other convenient choices are also appropriate; cf. [28]. The continuous LBB condition 41 holds thanks to the surjectivity of the divergence operator from $H(\mathrm{div},\Omega)$ to $L^{2}(\Omega)$; see, e.g., [25, Lemma 51.2]. To ensure that 40 holds, a natural choice of subspace would be the $H(\mathrm{div})$-conforming Raviart–Thomas space for $V_{h}$ and the broken polynomial space of the same order for $W_{h}$; see, e.g., [25, Lemma 51.10].
3.2. The PG method is well-defined
We prove that each nonlinear subproblem 10 has a unique solution.
Theorem 3.1 (Existence and uniqueness of solutions).
Assume we are in the setting outlined in Sections 1.3 and 3.1. Then for every $k\geq 1$, the nonlinear saddle point problem 10a-10b admits a unique solution pair $(u_{h}^{k},\psi_{h}^{k})\in V_{h}\times W_{h}$.
Proof.
We drop the superscript $k$ to simplify notation and start the proof with the uniqueness assertion. Indeed, if $(\hat{u}_{h},\hat{\psi}_{h})$ and $(u_{h},\psi_{h})$ solve 10, then, assuming $\hat{u}_{h}\neq u_{h}$ and using coercivity of $a$ and the strict monotonicity of $\nabla\mathcal{R}^{*}$, we obtain
$\nu\|\hat{u}_{h}-u_{h}\|_{V}^{2}\leq\alpha a(\hat{u}_{h}-u_{h},\hat{u}_{h}-u_{h})$ $=-b(\hat{u}_{h}-u_{h},\hat{\psi}_{h}-\psi_{h})$
$=-(\nabla\mathcal{R}^{*}(\hat{\psi}_{h})-\nabla\mathcal{R}^{*}(\psi_{h}),\hat{\psi}_{h}-\psi_{h})_{\Omega_{d}}$
$<0,$
which implies $\hat{u}_{h}=u_{h}$. Equation 10a then implies $b(v_{h},\hat{\psi}_{h}-\psi_{h})=0$ for all $v_{h}\in V_{h}$ and the assumed compatibility assumptions on $V_{h}$ and $W_{h}$ yield $\hat{\psi}_{h}=\psi_{h}$. Indeed, a direct consequence of 40 is that we can estimate
(48) $\|b(\cdot,w)\|_{V_{h}^{\prime}}\geq\beta_{0}\|w\|_{Q^{\prime}},$
for all $w\in W_{h}$. In other words, the map $w\mapsto b(\cdot,w)$ is injective with closed range between the spaces $(W_{h},\|\cdot\|_{Q^{\prime}})\to(V_{h}^{\prime},\|\cdot\|_{V_{h}^{\prime}})$.
To show the existence of solutions, we consider the following Lagrangian $\mathcal{L}:V_{h}\times W_{h}\rightarrow\mathbb{R}$:
(49) $\mathcal{L}(v,w)=\frac{\alpha}{2}a(v,v)-\alpha F(v)+b(v,w)-b(v,\psi_{h}^{k-1})-(\mathcal{R}^{*}(w),1)_{\Omega_{d}}.$
Clearly, every critical point of $\mathcal{L}$ is a solution to 10a-10b. We will show the existence of a critical point by minimizing in the first variable and maximizing in the second variable of $\mathcal{L}$. As $a$ is coercive, for fixed $w\in W_{h}$ we can find a unique $v(w)\in V_{h}$ satisfying
(50) $v(w)=\underset{v\in V_{h}}{\operatorname*{arg\,min}}\,\mathcal{L}(v,w).$
Note that $v(w)$ solves
(51) $\alpha[a(v(w),v)-F(v)]=b(v,\psi_{h}^{k-1})-b(v,w)~\text{ for all }v\in V_{h},$
or equivalently, setting $A=v\mapsto a(v,\cdot)$, we have
(52) $v(w)=A^{-1}[F+\alpha^{-1}b(\cdot,\psi_{h}^{k-1}-w)].$
Substituting $v=v(w)\in V_{h}$ into 49, we obtain
$J(w)\coloneqq\mathcal{L}(v(w),w)=-\frac{1}{2}a(v(w),v(w))-\int_{\Omega_{d}}\mathcal{R}^{*}(w)\operatorname{d}\!\mathcal{H}_{d}.$
Now, employing formula 52 we bound $\|v(w)\|_{V}$ from below:
$\|v(w)\|_{V}$ $=\|A^{-1}[F+\alpha^{-1}b(\cdot,\psi_{h}^{k-1}-w)]\|_{V}$
$\geq c\|F+\alpha^{-1}b(\cdot,\psi_{h}^{k-1}-w)\|_{V_{h}^{\prime}}$
$\geq c\left[\|b(\cdot,w)\|_{V_{h}^{\prime}}-\|F+b(\cdot,\psi_{h}^{k-1})\|_{V_{h}^{\prime}}\right]$
$\geq c\left[\beta_{0}\|w\|_{Q^{\prime}}-1\right]$
where, in the first step, we used that the isomorphism $A^{-1}:V_{h}^{\prime}\to V_{h}$ is injective with closed range. In the last estimate, we used 48. Together with the fact that we can lower bound $(\mathcal{R}^{*}(w),1)_{\Omega_{d}}$ by an affine linear function and the equivalence of norms in finite dimensional spaces, we deduce that $J(w)\to\infty$ as $\|w\|_{Q^{\prime}}\to\infty$. The map $w\mapsto v(w)$ is continuous as the first term is affine linear and $\mathcal{R}^{*}$ is continuous, hence we can guarantee the existence of a maximizer $w^{*}$ for $J$.
We now show that the pair $(v(w^{*}),w^{*})$ is the sought-after saddle point of $\mathcal{L}$. This follows from standard arguments; see e.g., [16, Exercise 7.16-4]. We provide some details for completeness. First observe that since $\mathcal{L}(v,\cdot)$ is concave, we have that
(53) $\theta\mathcal{L}(v(\delta_{\theta}),w)+(1-\theta)\mathcal{L}(v(\delta_{\theta}),w^{*})\leq\mathcal{L}(v(\delta_{\theta}),\delta_{\theta})=J(\delta_{\theta})\leq J(w^{*}),$
for any $\theta\in[0,1]$, $w\in W_{h}$ and $\delta_{\theta}=\theta w+(1-\theta)w^{*}$. Thus,
(54) $\theta\mathcal{L}(v(\delta_{\theta}),w)\leq J(w^{*})+(\theta-1)\mathcal{L}(v(\delta_{\theta}),w^{*})\leq\theta J(w^{*})=\theta\mathcal{L}(v(w^{*}),w^{*}),$
where we used that $\mathcal{L}(v(w^{*}),w^{*})\leq\mathcal{L}(v(\delta_{\theta}),w^{*})$, see 50. Then, we obtain that for any $\theta>0$, $\mathcal{L}(v(\delta_{\theta}),w)\leq\mathcal{L}(v(w^{*}),w^{*})$. With the continuity of the map $w\mapsto v(w)$ and of $\mathcal{L}(\cdot,w)$, we conclude that
$\mathcal{L}(v(w^{*}),w)\leq\mathcal{L}(v(w^{*}),w^{*})~\text{ for all }w\in W_{h}.$
Hence,
(55) $\inf_{v_{h}\in V_{h}}\sup_{w\in W_{h}}\mathcal{L}(v,w)\leq\sup_{w\in W_{h}}\mathcal{L}(v(w^{*}),w)\leq\mathcal{L}(v(w^{*}),w^{*})=\sup_{w\in W_{h}}\inf_{v\in V_{h}}\mathcal{L}(v,w).$
Since $\sup_{w\in W_{h}}\inf_{v\in V_{h}}\mathcal{L}(v,w)\leq\inf_{v_{h}\in V_{h}}\sup_{w\in W_{h}}\mathcal{L}(v,w)$ always holds, we conclude that
$\sup_{w\in W_{h}}\mathcal{L}(v(w^{*}),w)=\mathcal{L}(v(w^{*}),w^{*})=\inf_{v\in V_{h}}\mathcal{L}(v,w^{*}),$
which finishes the proof. ∎
In addition to the existence and uniqueness result of Theorem 3.1, we seek stability bounds on the iterates $(u_{h}^{k},\psi_{h}^{k})$, showing that these discrete solutions $u_{h}^{k}$ remain uniformly bounded in suitable norms independently of $h$, $\{\alpha_{k}\}$ and $\ell$. Uniform stability of $\psi_{h}^{\ell}$ in weak norms with respect to $h$ is also expected. Such bounds can be established under additional technical assumptions on $\nabla\mathcal{R}^{*}$, with details provided in Appendix A; in particular, see Theorem A.1.
3.3. Energy dissipation
We establish an energy dissipation property, which serves as a key tool for proving both stability and convergence of Algorithm 1.
Lemma 3.2 (Energy dissipation).
The following property holds for all $k\geq 1$:
(56) $E(u_{h}^{k+1})+\frac{1}{\alpha_{k+1}}(\mathcal{D}^{*}(\psi_{h}^{k+1},\psi_{h}^{k})+\mathcal{D}^{*}(\psi_{h}^{k},\psi_{h}^{k+1}),1)_{\Omega_{d}}\leq E(u_{h}^{k}).$
Proof.
Utilizing the coercivity of $a$ (8) and (10a)–(10b), we directly observe that
(57) $E(u_{h}^{k+1})$ $\leq E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k+1}),u_{h}^{k+1}-u_{h}^{k}\rangle$
$=E(u_{h}^{k})+\frac{1}{\alpha_{k+1}}b(u_{h}^{k+1}-u_{h}^{k},\psi_{h}^{k}-\psi_{h}^{k+1})$
$=E(u_{h}^{k})-\frac{1}{\alpha_{k+1}}(\nabla\mathcal{R}^{*}(\psi_{h}^{k+1})-\nabla\mathcal{R}^{*}(\psi_{h}^{k}),\psi_{h}^{k+1}-\psi_{h}^{k})_{\Omega_{d}}.$
The three points identity (36) provides the result. ∎
3.4. Best approximation error estimates
We now derive a priori best approximation estimates on the error between the discrete iterates $u_{h}^{k}$ of Algorithm 1 and the true solution $u^{*}$ of 4. In addition, we derive estimates between the dual variables $\lambda_{h}^{k}$ and $\lambda^{*}$ and the observables $o_{h}^{k}$ and $o^{*}$. These estimates yield convergence rates and a general mesh-independence property, even for low-regularity solutions.
Theorem 3.3 (Best approximation of the primal iterates $u_{h}^{\ell}$).
Let $u^{*}\in K$ be the solution to 4, let $\{u_{h}^{k}\}_{k=1}^{\ell}$ be defined via Algorithm 1. Assume we are in the setting outlined in Sections 1.3 and 3.1. Then the following estimate is valid for every $\ell\geq 1$:
(58) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{V}\\ +|\langle E^{\prime}(u^{*}),\Pi_{h}u^{*}-u^{*}\rangle|+\inf_{v\in K}|\langle E^{\prime}(u^{*}),v-u_{h}^{\ell}\rangle|,$
where we recall that $o^{*}=Bu^{*}$, $o_{h}^{0}=\nabla\mathcal{R}^{*}(\psi_{h}^{0})$, and $\Pi_{h}$ is the Fortin map given in (42).
Proof.
Observe that with the coercivity property of $a$ (8), we have for any $u,v\in V$
(59) $E(u)-E(v)=\langle E^{\prime}(v),u-v\rangle+\frac{1}{2}a(u-v,u-v)\geq\langle E^{\prime}(u),u-v\rangle+\frac{\nu}{2}\|u-v\|^{2}_{V},$
Setting $v=u_{h}^{k}$ above, we arrive at the following inequality for any $u\in V$:
$E(u)$ $\geq E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k}),u-u_{h}^{k}\rangle+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}$
$=E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u_{h}^{k}\rangle+\langle E^{\prime}(u_{h}^{k}),u-\Pi_{h}u\rangle+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}.$
We now rewrite the second term above using (10a), the definition of the dual variable $\lambda_{h}^{k}$ (13), and the Fortin map (42). We obtain
(60) $\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u_{h}^{k}\rangle=b(\Pi_{h}u-u_{h}^{k},\lambda_{h}^{k})=b(u-u_{h}^{k},\lambda_{h}^{k}).$
Proceeding, we use some ideas from [21, 27]. Namely, we use (10b), the definition of $o_{h}^{k}$ (12), the observation that $\psi_{h}^{k}=\nabla\mathcal{R}(o_{h}^{k})$ from (26), and the three points identity (36). This yields
(61) $b(u-u_{h}^{k},\lambda_{h}^{k})=(Bu-Bu_{h}^{k},\lambda_{h}^{k})_{\Omega_{d}}$ $=\frac{1}{\alpha_{k}}(Bu-o_{h}^{k},\nabla\mathcal{R}(o_{h}^{k-1})-\nabla\mathcal{R}(o_{h}^{k}))_{\Omega_{d}}$
$=\frac{1}{\alpha_{k}}(\mathcal{D}(Bu,o_{h}^{k})-\mathcal{D}(Bu,o_{h}^{k-1})+\mathcal{D}(o_{h}^{k},o_{h}^{k-1}),1)_{\Omega_{d}}.$
Collecting the above and using $\mathcal{D}(o_{h}^{k},o_{h}^{k-1})\geq 0$, we obtain the following identity:
(62) $E(u_{h}^{k})-E(u)+\frac{1}{\alpha_{k}}(\mathcal{D}(Bu,o_{h}^{k})-\mathcal{D}(Bu,o_{h}^{k-1}),1)_{\Omega_{d}}+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}\\ \leq\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u\rangle.$
Proceeding, we multiply by $\alpha^{k}$, sum from $k=1$ to $k=\ell$, and use the energy dissipation property Lemma 3.2. This yields
(63) $(E(u_{h}^{\ell})-E(u))\sum_{k=1}^{\ell}\alpha_{k}+\frac{\nu}{2}\sum_{k=1}^{\ell}\alpha_{k}\|u-u_{h}^{k}\|_{V}^{2}+(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}\\ \leq(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}+\sum_{k=1}^{\ell}\alpha_{k}\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u\rangle.$
We now define the weighted average $\overline{u}_{h}^{\ell}=\sum_{k=1}^{\ell}\alpha_{k}u_{h}^{k}/\sum_{k=1}^{\ell}\alpha_{k}$ and use Jensen’s inequality
(64) $\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}\leq\frac{\sum_{k=1}^{\ell}\alpha_{k}\|u-u_{h}^{k}\|_{V}^{2}}{\sum_{k=1}^{\ell}\alpha_{k}}.$
Thus, upon dividing 63 by $\sum_{k=1}^{\ell}\alpha_{k}$ then adding and subtracting $\langle E^{\prime}(u),\Pi_{h}u-u\rangle$, we obtain
(65) $E(u_{h}^{\ell})-E(u)+\frac{1}{2}\nu\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}+\frac{(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\\ \leq\frac{(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(\overline{u}_{h}^{\ell})-E^{\prime}(u),\Pi_{h}u-u\rangle+\langle E^{\prime}(u),\Pi_{h}u-u\rangle.$
Proceeding, we use the continuity of $a$ (8) and Young’s inequality to bound
$\langle E^{\prime}(\overline{u}_{h}^{\ell})-E^{\prime}(u),\Pi_{h}u-u\rangle\leq M\|\overline{u}_{h}^{\ell}-u\|_{V}\|\Pi_{h}u-u\|_{V}\leq\frac{\nu}{2}\|\overline{u}_{h}^{\ell}-u\|_{V}^{2}+\frac{M^{2}}{2\nu}\|\Pi_{h}u-u\|_{V}^{2}.$
Combining the above estimates yields
(66) $E(u_{h}^{\ell})-E(u)+\frac{(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\\ \leq\frac{(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u-u\|^{2}_{V}+\langle E^{\prime}(u),\Pi_{h}u-u\rangle.$
At this stage, we select $u=u^{*}$ and note that for any $v\in K$, 59 and 9 provide the bound
$E(u_{h}^{\ell})-E(u^{\ast})+\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ $\geq\langle E^{\prime}(u^{\ast}),v-u^{\ast}\rangle+\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\geq\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\,.$
Thus, adding $\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ to both sides of 66 and recalling that $Bu^{*}=o^{*}$, we arrive at
(67) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}+\frac{(\mathcal{D}(o^{*},o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{V}\\ +|\langle E^{\prime}(u^{*}),\Pi_{h}u^{*}-u^{*}\rangle|+|\langle E^{\prime}(u^{*}),v-u_{h}^{\ell}\rangle|.$
The result follows because the choice $v$ was arbitrary and $\mathcal{D}(o^{*},o_{h}^{\ell})\geq 0$. ∎
The next two results relate the error in $u_{h}^{\ell}$ to the errors in the dual variables $\lambda_{h}^{\ell}=(\psi_{h}^{\ell-1}-\psi_{h}^{\ell})/\alpha^{\ell}$ and observables $o_{h}^{\ell}=\nabla\mathcal{R}^{*}(\psi_{h}^{\ell})$.
Lemma 3.4 (Best approximation of the dual variables $\lambda_{h}^{\ell}$).
There exists a constant $\beta>0$, such that for any $\ell\geq 1$,
(68) $\beta\|\lambda^{*}-\lambda_{h}^{\ell}\|_{Q^{\prime}}\leq\sup_{v\in V}\frac{|\langle E^{\prime}(u^{*}),v-\Pi_{h}v\rangle|}{\|v\|_{V}}+M\|\Pi_{h}\|\|u^{*}-u_{h}^{\ell}\|_{V}.$
Proof.
We first estimate $\|B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell})\|_{V^{\prime}}$. For any $v\in V$, we use the Fortin operator 42 and equation 10a to write
$\langle B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell}),v\rangle$ $=\langle B^{\prime}\lambda^{*},v-\Pi_{h}v\rangle+\langle B^{\prime}\lambda^{*},\Pi_{h}v\rangle-\frac{1}{\alpha^{\ell}}b(v,\psi_{h}^{\ell-1}-\psi_{h}^{\ell})$
$=\langle E^{\prime}(u^{*}),v-\Pi_{h}v\rangle+\langle E^{\prime}(u^{*}),\Pi_{h}v\rangle-\frac{1}{\alpha^{\ell}}b(\Pi_{h}v,\psi_{h}^{\ell-1}-\psi_{h}^{\ell})$
$=\langle E^{\prime}(u^{*}),v-\Pi_{h}v\rangle+\langle E^{\prime}(u^{*})-E^{\prime}(u_{h}^{\ell}),\Pi_{h}v\rangle.$
Invoking the continuity of the bilinear form $a$ and the map $\Pi_{h}$, we obtain that
(69) $\|B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell})\|_{V^{\prime}}\leq\sup_{v\in V}\frac{|\langle E^{\prime}(u^{*}),v-\Pi_{h}v\rangle|}{\|v\|_{V}}+M\|\Pi_{h}\|\|u^{*}-u_{h}^{\ell}\|_{V}.$
Since the map $B:V\rightarrow Q$ is surjective, it follows from the closed range Theorem, see e.g. [22, Lemma A.40], that there exists a constant $\beta>0$ with
(70) $\beta\|\lambda^{*}-\lambda_{h}^{\ell}\|_{Q^{\prime}}\leq\|B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell})\|_{V^{\prime}}.$
Combining 69 with 70 finishes the proof. ∎
Lemma 3.5 (Best approximation of the discrete observables $o_{h}^{\ell}$).
The following estimate holds for all $\ell\geq 1$:
(71) $\|Bu^{*}-o_{h}^{\ell}\|_{L^{2}(\Omega_{d})}\leq\|B(u^{*}-u_{h}^{\ell})\|_{L^{2}(\Omega_{d})}+\|(I-P_{W_{h}})(Bu_{h}^{\ell}-o_{h}^{\ell})\|_{L^{2}(\Omega_{d})},$
where $P_{W_{h}}$ is the $L^{2}$-projection operator onto $W_{h}$ 75.
Proof.
This is a direct consequence of the triangle inequality and the observation that $P_{W_{h}}Bu_{h}^{\ell}=P_{W_{h}}o_{h}^{\ell}$, which follows from 10b. ∎
3.5. Convergence rates and mesh-independence
We now derive abstract error convergence rates from Theorem 3.3 and Lemmas 3.4 and 3.5. For simplicity, we focus on the following spaces $V\subset H^{1}(\Omega)$ and $V\subset H^{1}(\Omega;\mathbb{R}^{n})$. Hereafter, we do not explicitly differentiate between scalar and vector-valued spaces, using the notation for the scalar space while noting that the same results hold, mutatis mutandis, for vector-valued spaces. To derive error rates, we require the construction of an enriching map that allows us to control the last term in (58). In turn, we make the following general assumption, which is critical to defining this map. First, we motivate this assumption with a remark.
Remark 1 (Motivating Assumption 2).
Equation 10b allows the PG method to be viewed as a partially-nonconforming finite element method in the sense that the approximations $\{u_{h}^{\ell}\}$ generally do not belong to the feasible set $K$, where the true solution $u^{*}$ resides. Instead, 10b characterizes an approximate feasible set $K_{h}\not\subset K$ containing the iterates $\{u_{h}^{\ell}\}$. For example, consider the obstacle problem, Example 1.1, with the Shannon entropy from Example 2.1, and suppose that $W_{h}=\mathbb{P}_{0}(\mathcal{T}_{h})$. Then, by choosing $w_{h}=\chi_{T}$ (the indicator function of an element $T$), we see that
(72) $u_{h}^{\ell}\in K_{h}=\left\{v_{h}\in V_{h}\mid\int_{K}(v_{h}-\phi)>0\right\}.$
As expected from the analysis of nonconforming discretizations of obstacle problems [5, Section 5.2.1], one then requires an operator $\mathcal{E}_{h}:K_{h}\rightarrow K$ with suitable approximation properties. See Section 4.3 for an explicit construction of the enriching map for $K_{h}$ defined in 72.
Assumption 2.
There exists a continuous map $\mathcal{E}_{h}:V\rightarrow V$ with the property:
(73) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{E}_{h}:K\cup K_{h}\rightarrow K$},
where
(74) $K_{h}=\{u_{h}\in V_{h}\mid P_{W_{h}}(Bu_{h}-\nabla\mathcal{R}^{*}(\psi))=0\text{ for some }\psi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})\},$
and $P_{W_{h}}$ is the $L^{2}(\Omega_{d};\mathbb{R}^{m})$-projection operator onto $W_{h}$, defined by
(75) $(P_{W_{h}}o,w_{h})_{\Omega_{d}}=(o,w_{h})_{\Omega_{d}}~\text{ for all }o\in L^{2}(\Omega_{d};\mathbb{R}^{m}),~w_{h}\in W_{h}.$
The set $K_{h}$ contains functions $u_{h}\in V_{h}$ such that $Bu_{h}\in W_{h}$ is the $L^{2}$-projection of a constrained observable $o\in\mathcal{O}$. It is constructed to ensure that $u_{h}^{k}\in K_{h}$ for all iterations $k$; see 10b. To derive error rates, we require the following quasi-interpolation assumption on the stability and approximation of the Fortin and enriching operators. The decomposition of the enriching map given in 77 reflects the constructions used in Sections 4 and 5, below.
Assumption 3.
Assume that the Fortin operator satisfying 42 is stable in the sense that
(76a) $\|\Pi_{h}v\|_{L^{2}(\Omega)}$ $\lesssim\|v\|_{L^{2}(\Omega)}+h\|\nabla v\|_{L^{2}(\Omega)},$
(76b) $\|\nabla(\Pi_{h}v)\|_{L^{2}(\Omega)}$ $\lesssim\|\nabla v\|_{L^{2}(\Omega)},$
for all $v\in V$. Further, assume that the enriching map $\mathcal{E}_{h}$ of Assumption 2 is affine linear:
(77) $\mathcal{E}_{h}w=\mathcal{C}_{h}w+\varepsilon,$
where $\mathcal{C}_{h}:V\rightarrow V_{h}$ is linear and $\varepsilon\in V$.
For $0\leq s\leq 1$, $0\leq t\leq 1$, and $w\in H^{1+s}(\Omega)\cap V$, assume that
(78a) $\|w-\Pi_{h}w\|_{L^{2}(\Omega)}+h\|\nabla(w-\Pi_{h}w)\|_{L^{2}(\Omega)}$ $\lesssim h^{1+s}|w|_{H^{1+s}(\Omega)},$
(78b) $\|w-\mathcal{C}_{h}w\|_{L^{2}(\Omega)}+h\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(\Omega)}$ $\lesssim h^{1+s}|w|_{H^{1+s}(\Omega)},$
(78c) $\|\varepsilon\|_{H^{t}(\Omega)}$ $\lesssim h^{1+s-t}.$
Theorem 3.6 (Convergence rates).
Assume that $u^{*}\in H^{1+s}(\Omega)$ and $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$ for some $s,r\in(0,1]$, and let $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega_{d};\mathbb{R}^{m})}\lesssim 1$. Let Assumptions 1, 2 and 3 hold. Then
(79) $\|u^{*}-u_{h}^{\ell}\|_{H^{1}(\Omega)}^{2}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{Q^{\prime}}\lesssim\frac{C_{\mathrm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}\,h^{2\cdot\min\{r,s\}}$
for all $\ell\geq 1$ and $h>0$, where $C_{\mathrm{stab}}$ and $C_{\mathrm{reg}}$ are positive constants independent of $h$, $\ell$, and the parameters $\alpha_{k}$.
Proof.
Since $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$, we further bound the last two terms in 58 by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}$ and by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$ respectively. Thus, we obtain
(80) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{H^{1}(\Omega)}\\ +\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}+\inf_{v\in K}\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}.$
We now proceed by bounding each term in 80. Recalling that $o_{h}^{0}=\nabla\mathcal{R}^{*}(\psi_{h}^{0})$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega_{d};\mathbb{R}^{m})}\lesssim 1$, and the definition of $\mathcal{D}$ (33), it readily follows that
(81) $\mathcal{D}(o^{*},o_{h}^{0})\leq C_{\mathrm{stab}},$
for some constant $C_{\mathrm{stab}}$ independent of $h,\ell$ and $\alpha_{k}$ since $\mathcal{R}$ is continuous on $\mathcal{O}$ and $\nabla\mathcal{R}^{*}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$.
Proceeding, Assumption 3 directly provides the bound
(82) $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1}(\Omega)}^{2}\lesssim h^{2s}|u^{*}|^{2}_{H^{1+s}(\Omega)}.$
To handle the $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1-r}(\Omega)}$ term in 80, we first note that from space interpolation between $L^{2}(\Omega)$ and $H^{1}(\Omega)$ (see, e.g., [8, Chapter 14, Proposition 14.1.5]) and 78a, we obtain that for any $w\in H^{1+s}(\Omega)$,
(83) $\|w-\Pi_{h}w\|_{H^{t}(\Omega)}+\|w-\mathcal{C}_{h}w\|_{H^{t}(\Omega)}\lesssim h^{1+s-t}|w|_{H^{1+s}(\Omega)},\quad s,t\in[0,1].$
Therefore, we obtain that
(84) $\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}\lesssim h^{r+s}|u^{*}|_{H^{1+s}(\Omega)}.$
It remains to estimate the term in 80 involving $\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$. To this end, we choose $v=\mathcal{E}_{h}(u_{h}^{\ell})$. This is a valid choice since $u_{h}^{\ell}\in K_{h}$ by 10b and $\mathcal{E}_{h}u_{h}^{\ell}\in K$ by Assumption 2. With the help of 78 and (83), we now formulate the following estimate:
(85) $\|u_{h}^{\ell}-\mathcal{E}_{h}u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$ $=\|(u_{h}^{\ell}-u^{*})-\mathcal{C}_{h}(u_{h}^{\ell}-u^{*})+(u^{*}-\mathcal{E}_{h}u^{*})\|_{H^{1-r}(\Omega)}$
$\leq\|(u_{h}^{\ell}-u^{*})-\mathcal{C}_{h}(u_{h}^{\ell}-u^{*})\|_{H^{1-r}(\Omega)}+\|u^{*}-\mathcal{E}_{h}u^{*}\|_{H^{1-r}(\Omega)}$
$\leq ch^{r}\|\nabla(u_{h}^{\ell}-u^{*})\|_{L^{2}(\Omega)}+\tilde{c}h^{r+s}(|u^{*}|_{H^{1+s}(\Omega)}+1)$
$\leq\frac{1}{4}\nu\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}^{-1}\|\nabla(u_{h}^{\ell}-u^{*})\|_{L^{2}(\Omega)}^{2}+c^{2}\nu^{-1}\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}h^{2r}$
$+\tilde{c}h^{r+s}(|u^{*}|_{H^{1+s}(\Omega)}+1),$
where $c$ and $\tilde{c}$ are mesh-independent constants. Incorporating 81, 82, 84, and 85 into 80 yields the required bound on $\|u_{h}^{\ell}-u^{*}\|_{H^{1}(\Omega)}$.
The estimate on $\|\lambda^{*}-\lambda_{h}^{\ell}\|_{H^{-1}(\Omega)}$ follows from Lemma 3.4. In particular, the first term in 68 is bounded by:
$\sup_{v\in H^{1}(\Omega)}\frac{|\langle E^{\prime}(u^{*}),v-\Pi_{h}v\rangle|}{\|v\|_{H^{1}(\Omega)}}\leq\sup_{v\in H^{1}(\Omega)}\frac{\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-\Pi_{h}v\|_{H^{1-r}(\Omega)}}{\|v\|_{H^{1}(\Omega)}}\lesssim h^{r},$
using 83 for the second inequality. ∎
The final result of this section follows from the fact that the approximation error in 79 is controlled by independent optimization and discretization error terms, each depending only on $\ell$ and $h$, respectively.
Corollary 1 (Asymptotic mesh-independence).
Let $\epsilon>0$. Under the assumptions of Theorem 3.6, there exists a critical mesh size $h_{\epsilon}>0$ and iteration number $\ell_{\epsilon}\geq 1$ such that
(86) $\|u^{\ast}-u^{\ell}_{h}\|_{V}+\|\lambda^{\ast}-\lambda^{\ell}_{h}\|_{Q^{\prime}}\leq\epsilon~\text{ for all }0<h\leq h_{\epsilon}\text{ and }\ell\geq\ell_{\epsilon}\,.$
3.6. Discussion of results
Theorems 3.3 and 3.6 rigorously demonstrate three important features of the PG method, Algorithm 1, for the first time. To date, these features have only been observed numerically [36, 28, 19, 40].
  • First, the PG iterates $u^{\ell}_{h}$ converge to the true solution $u^{\ast}$ as $\ell\to\infty$ and $h\to 0$, even with bounded proximity parameters, $\alpha_{k}\leq\textrm{const.}$ This stands in contrast to penalty [30, Chapter 1.7] and interior point methods [38, Chapter 19] — and even augmented Lagrangian methods in infinite-dimensional spaces [35, 3] — which all require taking a relaxation parameter to a singular limit.
  • Second, the number of iterations is asymptotically independent of the mesh size. More specifically, Corollary 1 shows that a user can guarantee convergence to any desired accuracy by independently selecting the mesh size $h$ and iteration count $\ell$. Such mesh-independence is an important feature [45] seen in, e.g., the primal-dual active set method [32], but for a more restrictive class of problems than considered here [34]. Moreover, note that Corollary 1 is a global result, holding for any uniformly bounded initial guess $\psi_{h}^{0}$. In contrast, the celebrated mesh-independence of Newton’s method [1, 50] is only a local property, as it requires an accurate initial guess on a sufficiently fine mesh.
  • Finally, PG can be applied without modification to problems with low-regularity solutions or multipliers. Indeed, Theorem 3.6 provides sufficient conditions on $\Pi_{h}$ and $\mathcal{E}_{h}$ to obtain optimal convergence rates in $h$ depending naturally on the solution $u^{\ast}$ and Fréchet Derivative $E^{\prime}(u^{\ast})$. The following two sections show that these abstract conditions are checkable in practice. It is well-known that penalty methods can also be applied without modification in the low-regularity setting [33]. However, to the best of our knowledge, the existing theory for discretized penalty methods requires $u^{*}\in H^{1+s}(\Omega)$ and $s\in(1/2,1]$ when applied to the obstacle and Signorini problems. See [44, 31] for the obstacle problem and [14, Section 6] for the Signorini problem. Notably, the PG results hold for any $s\in(0,1]$.
Remark 2 (Higher-order convergence rates).
The current estimates deliver at most first-order convergence rates, in which case it is required that $u^{\ast}\in H^{2}(\Omega)$ and $E^{\prime}(u^{*})\in L^{2}(\Omega)$. However, the PG method is not limited to low-degree polynomial subspaces $V_{h}\times W_{h}$ and, notably, high-order rates have been numerically observed in several studies with high-degree subspaces [36, 28, 40]. Thus, a crucial question remains unanswered: If the exact solution is sufficiently smooth, then what are the most general conditions that guarantee high-order convergence rates? Indeed, a major difficulty in extending the current framework to the examples below lies in constructing high-order positivity-preserving approximations, which are encoded in the map $\mathcal{E}_{h}$. This is known to be a challenging task, impeded by some impossibility results [39].
4. Application I: The obstacle problem
In this section, we apply the general framework developed in Section 3 to derive error estimates for the following unilateral obstacle problem, see also Example 1.1:
(87) $\min_{v\in K}E(v),\quad E(v)=\frac{1}{2}\|\nabla v\|^{2}_{L^{2}(\Omega)}-(f,v)_{\Omega},$
where the closed and convex set $K$ is
(88) $K=\{v\in H^{1}_{0}(\Omega)\mid v\geq\phi\text{ a.e.\ in }\Omega\},$
and the obstacle $\phi\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ is fixed. For simplicity, we assume that $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi|_{\partial\Omega}=-\delta\leq 0$} for some non–negative constant $\delta$. It is well-known that there exists a unique solution to 87 [16, Theorem 6.9-1], here denoted by $u^{\ast}$.
For completeness, we write the PG method for the obstacle problem in Algorithm 2. In this case, we have used
(89) $\mathcal{R}(u)=(u-\phi)\ln(u-\phi)-(u-\phi),$
which implies that
(90) $\nabla\mathcal{R^{*}}(\psi)=\exp(\psi)+\phi.$
Assumption 1 holds for this choice of Legendre function [36, Theorem 4.1]; see also [36, Proposition A.8]. Alternative choices for $\mathcal{R}$ are also admissible, without affecting the proceeding error analysis.
Algorithm 2 The Proximal Galerkin Method for the Obstacle Problem
1:input: Initial latent solution guess $\psi_{h}^{0}\in W_{h}$, a sequence of positive proximity parameters $\{\alpha_{k}\}$.
2:Initialize $k=1$.
3:repeat
4: Find $u_{h}^{k}\in V_{h}$ and $\psi_{h}^{k}\in W_{h}$ such that
(91a) $\alpha_{k}\,(\nabla u^{k}_{h},\nabla v_{h})+(v_{h},\psi^{k}_{h}-\psi_{h}^{k-1})$ $=\alpha_{k}(f,v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(91b) $(u^{k}_{h},w_{h})-(\exp(\psi^{k}_{h})+\phi,w_{h})$ $=0$  $~\text{ for all }w_{h}\in W_{h}.$
5: Assign $k\leftarrow k+1$.
6:until a convergence test is satisfied.
In this section, where $\Omega_{d}=\Omega$ and $B=\mathrm{id}$, we use the following notation to remain consistent with [36], where the PG method was first introduced:
(92) $\tilde{u}_{h}^{k}:=o_{h}^{k}=\exp(\psi_{h}^{k})+\phi.$
We also consider two different choices for the finite element subspaces $V_{h}\times W_{h}$, both also introduced in [36].
Main goal: We derive error estimates in Corollary 2 for the PG method applied to the obstacle problem (Algorithm 2) for the two choices of $V_{h}\times W_{h}$ given in 94 and 95, below. To this end, we utilize Theorem 3.6 and verify that Assumptions 2 and 3 hold for each pair of spaces, respectively. We follow with preliminary results for each case.
Case I. $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ Define the space
(93) $\mathbb{B}(T)=\operatorname{span}\{b_{T}\},$
where the bubble function $b_{T}\colon T\to\mathbb{R}$ is the product of the linear nodal basis functions of the element $T$. The pair $(V_{h},W_{h})$ is then defined as
(94) $V_{h}$ $=\{v\in L^{\infty}(\Omega)\mid v|_{T}\in\mathbb{P}_{1}(T)\oplus\mathbb{B}(T)~\text{ for all }T\in\mathcal{T}_{h}\}\cap H^{1}_{0}(\Omega),$
$W_{h}$ $=\mathbb{P}_{0}(\mathcal{T}_{h}).$
Case II. $(\mathbb{P}_{1},\mathbb{P}_{1})$, i.e., continuous Lagrange elements. Define
(95) $V_{h}=W_{h}=\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega).$
Refer to Remark 3 for enforcing Dirichlet boundary conditions with these elements.
Remark 3 (Dirichlet boundary conditions).
Both of the $V_{h}$ subspaces in 94 and 95 imply fixing the degrees of freedom of $u_{h}^{k}$ on the Dirichlet boundary. This is a textbook procedure that is simplified in our setting because of the homogeneous boundary conditions specified in 87. Likewise, the continuous Lagrange elements defining $W_{h}$ in 95 also require fixing the degrees of freedom of the latent variable $\psi_{h}^{k}$ on the Dirichlet boundary. In this case, we construct the lift
$\psi_{0,h}:=\sum_{z\in\mathcal{N}_{h}\cap\partial\Omega}\nabla\mathcal{R}(0)(z)\varphi_{z},$
where $\mathcal{N}_{h}\cap\partial\Omega$ are the boundary dofs and $\{\varphi_{z}\}_{z\in\mathcal{N}_{h}}$ are the global shape functions corresponding to the (nodal) dofs $\mathcal{N}_{h}$. This construction ensures that the bound-preserving discrete solution $\tilde{u}_{h}^{k}$ 92 satisfies the same homogeneous Dirichlet boundary conditions as the solution $u^{\ast}$ it’s meant to approximate. Indeed, observe that
$\tilde{u}_{h}^{k}|_{\partial\Omega}=\nabla\mathcal{R}^{*}(\psi_{h}^{k}|_{\partial\Omega}+\psi_{0,h}|_{\partial\Omega})|_{\partial\Omega}=\nabla\mathcal{R}^{*}(\nabla\mathcal{R}(0))|_{\partial\Omega}=0,$
with $\nabla\mathcal{R}^{*}(\psi)=\exp(\psi)+\phi$ and $\nabla\mathcal{R}(u)=\ln(u-\phi)$ for the particular choice of Legendre function in Algorithm 2. In turn, when using the elements in Case II, we understand 91b as
(96) $(u_{h}^{k},w_{h})-(\nabla\mathcal{R}^{*}(\psi_{h}^{k}+\psi_{0,h}),w_{h})=0~\text{ for all }w_{h}\in W_{h}.$
4.1. Preliminaries for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$-element pair
We begin with the Fortin operator for the subspaces in 94, and establish the operator’s approximation properties required in Assumption 3.
Lemma 4.1 (Fortin operator).
The subspaces given in 94 satisfies the inf-sup condition 40 and there exists a stable Fortin operator satisfying 42 and
(97) $|\Pi_{h}w|_{H^{m}(\Omega)}\lesssim|w|_{H^{m}(\Omega)}\quad~\text{ for all }w\in H^{m}(\Omega),\;m\in\{0,1\}.$
In addition, for every $0\leq s\leq 1$ and for all $w\in H^{1+s}(\Omega)\cap H^{1}_{0}(\Omega),$ it holds that
(98) $\|w-\Pi_{h}w\|_{L^{2}(\Omega)}+h\|\nabla(w-\Pi_{h}w)\|_{L^{2}(\Omega)}$ $\lesssim h^{1+s}|w|_{H^{1+s}(\Omega)}.$
Proof.
The proof is based on the arguments in [36, Appendix B]. The operator $\Pi_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is constructed as follows:
(99) $\Pi_{h}=\tilde{\mathcal{I}}_{h}+\tilde{\Pi}_{h}(I-\tilde{\mathcal{I}}_{h}),$
where $\tilde{\mathcal{I}}_{h}:L^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega)$ is the quasi-interpolant introduced in [23, Section 6], $I$ is the identity operator, and $\tilde{\Pi}_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is defined element-wise to satisfy $(\tilde{\Pi}_{h}v)|_{T}:=b_{T}v_{T}\in\mathbb{B}(T)$, where $v_{T}\in\mathbb{P}_{0}(T)=\mathbb{R}$ solves
(100) $(b_{T}v_{T},\varphi)_{T}=(v,\varphi)_{T}~\text{ for all }\varphi\in\mathbb{P}_{0}(T).$
It is easy to see that $v_{T}=(v,1)_{T}/(b_{T},1)_{T}$ is the unique solution to the above. Owing to mesh regularity and to the properties of the bubble function [49, Lemma 4.1], we have
(101) $\|v_{T}\|^{2}_{L^{2}(T)}\lesssim(b_{T}v_{T},v_{T})_{T}=(v,v_{T})_{T}\leq\|v\|_{L^{2}(T)}\|v_{T}\|_{L^{2}(T)}.$
Therefore, we obtain
(102) $\|\tilde{\Pi}_{h}v\|_{L^{2}(T)}=\|b_{T}v_{T}\|_{L^{2}(T)}\leq\|v_{T}\|_{L^{2}(T)}\lesssim\|v\|_{L^{2}(T)}.$
Using 102 in 99 along with the triangle inequality and the stability of $\tilde{\mathcal{I}}_{h}$, we obtain the stability of $\Pi_{h}$ in $L^{2}(\Omega)$. To show stability in the $H^{1}$-semi norm, we apply triangle inequality, a local inverse estimate, 102 and the stability and approximation properties of $\tilde{\mathcal{I}}_{h}$ [23, Lemma 6.3 and Theorem 6.4]:
(103) $|\Pi_{h}w|_{H^{1}(T)}\lesssim|\tilde{\mathcal{I}}_{h}w|_{H^{1}(T)}+h_{T}^{-1}\|\tilde{\Pi}_{h}(I-\tilde{\mathcal{I}}_{h})w\|_{L^{2}(T)}\\ \lesssim|\tilde{\mathcal{I}}_{h}w|_{H^{1}(T)}+h_{T}^{-1}\|w-\tilde{\mathcal{I}}_{h}w\|_{L^{2}(T)}\lesssim|w|_{H^{1}(\Delta_{T})},$
where $\Delta_{T}$ is a macro-element. Summing over mesh elements and using mesh regularity shows stability in the $H^{1}(\Omega)$-seminorm. We conclude that 97 holds. This stability estimate and the observation that
$(\Pi_{h}v,w_{h})=(\tilde{\mathcal{I}}_{h}v,w_{h})+(\tilde{\Pi}_{h}(v-\tilde{\mathcal{I}}_{h}v),w_{h})=(v,w_{h})~\text{ for all }w_{h}\in W_{h}$
[…]
Note that the discrete iterates $u_{h}^{k}$ have bound-preserving local averages. Indeed, testing 91b by the indicator function of one element (an admissible test function because $W_{h}=\mathbb{P}_{0}(\mathcal{T}_{h})$) and using that $\nabla\mathcal{R}^{*}(\psi)>\phi$ shows that $u_{h}^{k}\in K_{h}$ for all $k$, with
(104) $K_{h}\subset\left\{v_{h}\in V_{h}\mid\int_{T}(v_{h}-\phi)\operatorname{d}\!x\geq 0\;~\text{ for all }T\in\mathcal{T}_{h}\right\}.$
To satisfy Assumption 2, we first require a map defined over $V$ and mapping the set $K_{h}$ to the set $K$. We will later shift this map to construct the enriching map $\mathcal{E}_{h}$, see 121. The natural choice is the Clément interpolant [17, 5]. Here, we utilize a specialized variant of this interpolant that employs weighted local averages over vertex patches to maintain second-order accuracy for smooth functions. This specific interpolant was first introduced in [29], but we slightly modify its definition to incorporate non-homogeneous boundary data.
Denote the centroid of an element $T\in\mathcal{T}_{h}$ by $s_{T}$; i.e., $s_{T}=(n+1)^{-1}\sum_{v\in\mathcal{V}_{T}}v,$ where $\mathcal{V}_{T}$ is the set of $n+1$ vertices of the element $T$. Since every node $z\in\mathcal{N}_{h}\backslash\partial\Omega$ belongs to the convex hull of the set $\{s_{T}\mid T\subset\omega_{z}\}$, we can write every $z\in\mathcal{N}_{h}\backslash\partial\Omega$ as a convex combination of local centroids:
(105) $z=\sum_{T\subset\omega_{z}}\alpha_{z,T}s_{T},\quad\sum_{T\subset\omega_{z}}\alpha_{z,T}=1,\quad\alpha_{z,T}\geq 0.$
Note that the choice of $\{\alpha_{z,T}\}_{T\subset\omega_{z}}$ is not unique if there are more than $n+1$ elements in $\omega_{z}$. We now define $\mathcal{C}_{h}:H^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}(\Omega)$ as follows:
(106) $\mathcal{C}_{h}v=\sum_{z\in\mathcal{N}_{h}}v_{z}\varphi_{z},\quad v_{z}=\sum_{T\subset\omega_{z}}\frac{\alpha_{z,T}}{|T|}\int_{T}v\operatorname{d}\!x\,\text{ if }z\in\mathcal{N}_{h}\backslash\partial\Omega.$
For $z\in\partial\Omega$, we set $v_{z}=(\mathcal{SZ}_{h}v)(z)$ where $\mathcal{SZ}_{h}:H^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})$ is the canonical Scott–Zhang interpolant defined in [46]. Note that with this definition, we recover that $(\mathcal{C}_{h}v-v)|_{\partial\Omega}=0$ for any $v$ with piecewise-polynomial boundary data.
Lemma 4.2 (Approximation properties of the modified Clément interpolant).
For any $T\in\mathcal{T}_{h}$ and for $0\leq s\leq 1$, it holds that
(107) $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}+h_{T}\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(T)}$ $\lesssim h_{T}^{1+s}|w|_{H^{1+s}(\Delta_{T})}$  $~\text{ for all }w\in H^{1+s}(\Delta_{T}),$
where $\Delta_{T}$ is the macro-element given by $\Delta_{T}=\cup_{z\in\mathcal{V}_{T}}\omega_{z}$.
Proof.
The following is a modified proof of [29, Theorem 11] for the case of inhomogeneous boundary data and $H^{1+s}(\Omega)$ functions. The key observation, as we show next, is that this interpolant preserves linear functions locally [29]; i.e., $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$ and $z\in\mathcal{N}_{h}$. For an interior node $z\in\mathcal{N}_{h}\backslash\partial\Omega$, it is clear that $b_{z}=b$ for a constant $b$ since $\sum_{T\subset\omega_{z}}\alpha_{z,T}=1$ and for a linear polynomial $p$, it follows that:
(108) $\mathcal{C}_{h}p(z)=\sum_{T\subset\omega_{z}}\alpha_{z,T}p(s_{T})=p\left(\sum_{T\subset\omega_{z}}\alpha_{z,T}s_{T}\right)=p(z),$
where we used that $\int_{T}p\operatorname{d}\!x=|T|p(s_{T})$ to arrive at the first equality. For a boundary node $z\in\mathcal{N}_{h}\cap{\partial\Omega}$, we conclude that $\mathcal{C}_{h}p(z)=(\mathcal{SZ}_{h}p)(z)=p(z)$. Therefore, $\mathcal{C}_{h}p=p$ for all $p\in\mathbb{P}_{1}(\Delta_{T})$ with $T\in\mathcal{T}_{h}$. This implies that
(109) $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}\leq\|w-p\|_{L^{2}(T)}+\|\mathcal{C}_{h}(p-w)\|_{L^{2}(T)}$
for any $T\in\mathcal{T}_{h}$ and $p\in\mathbb{P}_{1}(\Delta_{T})$. For the second term, we proceed by bounding $C_{h}(p-w)(z)$ for each $z\in\mathcal{V}_{T}.$ For an interior node $z\in\mathcal{N}_{h}\backslash\partial\Omega$, we apply Cauchy–Schwarz inequality and the fact that $0\leq\alpha_{z,T}\leq 1$:
(110) $|\mathcal{C}_{h}(p-w)(z)|\leq\sum_{T\subset\omega_{z}}|\alpha_{z,T}||T|^{-1}\|1\|_{L^{2}(T)}\|p-w\|_{L^{2}(T)}\lesssim|T|^{-1/2}\|p-w\|_{L^{2}(\omega_{z})}.$
In the above, we also used that $\mathrm{card}(\omega_{z})$ is uniformly bounded w.r.t. $h$ for all $z\in\mathcal{N}_{h}$, which follows from the shape-regularity of the mesh; see, e.g., [24, Proposition 11.6]. If $z\in\partial\Omega$, then $z\in\partial T\cap\omega_{z}$ for some $T\subset\omega_{z}$. Using the definition of $\mathcal{C}_{h}$ on boundary nodes, we apply local inverse and trace inequalities to bound
(111) $|\mathcal{C}_{h}(p-w)(z)|=|\mathcal{SZ}_{h}(p-w)(z)|$ $\lesssim h_{T}^{(1-n)/2}\|\mathcal{SZ}_{h}(w-p)\|_{L^{2}(\partial T)}$
$\lesssim h_{T}^{-n/2}\|\mathcal{SZ}_{h}(w-p)\|_{L^{2}(T)}$
$\lesssim h_{T}^{-n/2}(\|p-w\|_{L^{2}(\Delta_{T})}+h_{T}\|\nabla(p-w)\|_{L^{2}(\Delta_{T})}).$
In the last line above, we used the local stability property of $\mathcal{SZ}_{h}$ [46, Theorem 3.1]. Using that $C_{h}(p-w)|_{T}=\sum_{z\in\mathcal{V}_{T}}C_{h}(p-w)(z)\phi_{z}$ and that $\|\phi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$, we arrive at
(112) $\|\mathcal{C}_{h}(p-w)\|_{L^{2}(T)}\lesssim\|p-w\|_{L^{2}(\Delta_{T})}+h_{T}\|\nabla(p-w)\|_{L^{2}(\Delta_{T})}.$
Combining 109 with 112 and using the Bramble–Hilbert Lemma (see, e.g., [23, Lemma 5.6]) yields the required bound on the first term of 107. To obtain the required bound on the second term, we apply triangle and inverse estimates:
$\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(T)}\lesssim\|\nabla(w-\mathcal{SZ}_{h}w)\|_{L^{2}(T)}+h_{T}^{-1}(\|\mathcal{SZ}_{h}w-w\|_{L^{2}(T)}+\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}).$
The proof is completed by using the approximation properties of $\mathcal{SZ}_{h}$ [46, Theorem 4.1] and the proven bound on $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}$. ∎
4.2. Preliminaries for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair
For this case, the Fortin operator is simply the $L^{2}(\Omega)$-projection; cf. [36, Remark 5.2]. Thus, the main task is to construct the enriching map of Assumption 2, which will again depend on a modified Clément quasi-interpolant. Using the definition of $K_{h}$ 74 and testing with $\varphi_{z}$, we observe that
(113) $K_{h}\subset\left\{v_{h}\in V_{h}\mid\int_{\omega_{z}}(v_{h}-\phi)\varphi_{z}\operatorname{d}\!x\geq 0~\text{ for all }z\in\mathcal{N}_{h}\backslash\partial\Omega\right\},$
since the support of $\varphi_{z}$ is $\omega_{z}$, $\nabla\mathcal{R}^{*}(\psi_{h}^{\ell})-\phi>0$, and $\varphi_{z}\geq 0$. In what follows, we consider the quasi-interpolant proposed in [51, 12]:
(114) ${\mathcal{C}}_{h}v=\sum_{z\in\mathcal{N}_{h}}v_{z}\varphi_{z},\quad v_{z}=\frac{1}{\int_{\omega_{z}}\varphi_{z}}\int_{\omega_{z}}v\varphi_{z}\operatorname{d}\!x\text{ if }z\in\mathcal{N}_{h}\backslash\partial\Omega.$
If $z\in\partial\Omega$, we proceed as before and select $v_{z}=(\mathcal{SZ}_{h}v)(z)$. For simplicity, we make the following symmetry assumption on the mesh, which is sufficient for optimality of the quasi-interpolant given in 114. Note that this assumption is the same condition necessary for optimality of the classical Clément quasi-interpolant [29].
Assumption 4 (Local mesh symmetry).
For all $z\in\mathcal{N}_{h}\backslash\partial\Omega$, assume that
(115) $\frac{1}{|\omega_{z}|}\sum_{T\subset\omega_{z}}|T|s_{T}=z,\text{ where }s_{T}=(n+1)^{-1}\sum_{v\in\mathcal{V}_{T}}v.$
We now prove that this condition implies optimality of 114.
Lemma 4.3.
Suppose that the mesh $\mathcal{T}_{h}$ satisfies Assumption 4. For any $T\in\mathcal{T}_{h}$ and for $0\leq s\leq 1$,
(116) $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}+h_{T}\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(T)}$ $\lesssim h_{T}^{1+s}|w|_{H^{1+s}(\Delta_{T})}$  $~\text{ for all }w\in H^{1+s}(\Delta_{T}),$
where $\Delta_{T}$ is the macro-element given by $\Delta_{T}=\cup_{z\in\mathcal{V}_{T}}\omega_{z}$.
Proof.
The proof follows similar arguments to the proof of Lemma 4.2. We only highlight the key points. We first show that under 115, $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$. To see this first note that for linear $p$, we have
$\int_{T}p(x)\varphi_{z}(x)\operatorname{d}\!x=p\left(\sum_{z^{\prime}\in\mathcal{V}_{T}}\int_{T}z^{\prime}\varphi_{z^{\prime}}(x)\varphi_{z}(x)\operatorname{d}\!x\right)=\frac{|T|}{(n+1)(n+2)}p(z+(n+1)s_{T}),$
where we used that $\int_{T}\varphi_{z^{\prime}}\varphi_{z}\operatorname{d}\!x=(n+1)^{-1}(n+2)^{-1}|T|(1+\delta_{z^{\prime},z})$ [15, Exercise 4.1.1]. Hence,
(117) $\frac{1}{\int_{\omega_{z}}\varphi_{z}}\int_{\omega_{z}}p\varphi_{z}\operatorname{d}\!x=\frac{1}{(n+2)|\omega_{z}|}p\left(\sum_{T\subset\omega_{z}}|T|(z+(n+1)s_{T})\right)=p(z).$
In the last step, we used that $\int_{\omega_{z}}\varphi_{z}=|\omega_{z}|/(n+1)$ and 115. Along with the observation that $\overline{b}_{z}=b$ for any constant $b$, we conclude that $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$. In turn, it suffices to bound $\|\mathcal{C}_{h}(p-w)\|_{L^{2}(T)}$ as done in 109. If $z\in\mathcal{N}_{h}\backslash\partial\Omega$, we use the Cauchy–Schwarz inequality and the fact that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$ to deduce
(118) $|\mathcal{C}_{h}(p-w)(z)|\leq\sum_{T\subset\omega_{z}}(n+1)|\omega_{z}|^{-1}\|\phi_{z}\|_{L^{2}(T)}\|p-w\|_{L^{2}(T)}\lesssim h_{T}^{-n/2}\|p-w\|_{L^{2}(\omega_{z})}.$
The case $z\in\partial\Omega$ and the remaining steps follow identically to the proof of Lemma 4.2. The details are omitted for brevity. ∎
Remark 4 (Removing Assumption 4).
As suggested by the proof of Lemma 4.3, one can not guarantee that $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$ if 115 does not hold. Thus, a lack of local symmetry results in the proposed quasi-interpolant lacking optimality. Of course, one may be inspired by the construction of $\mathcal{C}_{h}$ in Section 4.1 to generalize the proposed operator 114. In particular, one can define a convex combination of weights $\{\alpha_{z^{\prime}}\}_{z^{\prime}\in\omega_{z}}$ such that the following reweighted operator,
$\mathcal{C}_{h}v(z)=\sum_{z^{\prime}\in\omega_{z}}\frac{\alpha_{z^{\prime}}}{\int_{\omega_{z^{\prime}}}\varphi_{z^{\prime}}}\int_{\omega_{z^{\prime}}}v\varphi_{z^{\prime}}\operatorname{d}\!x,\quad z\in\mathcal{N}_{h}\backslash\partial\Omega,$
preserves linear functions on the super-macro element $\cup_{z^{\prime}\in\mathcal{N}_{h}\cap\omega_{z}}\omega_{z^{\prime}}$. This construction requires a more delicate analysis and specific assumptions on $\mathcal{T}_{h}$ near the boundary $\partial\Omega$ to guarantee non-negativity of $\mathcal{C}_{h}(v_{h}-\phi)$ for all $v_{h}\in K_{h}\cup K$. Although we save the technical details for future work, we perform this type of construction on a one-dimensional subdomain for the Signorini problem considered in Section 5; in particular, see 137.
4.3. Error estimates for the obstacle problem
The preliminary results in Sections 4.1 and 4.2 leave us ready to state and prove optimal a priori error estimates for the PG method applied to the obstacle problem.
Corollary 2 (A priori error estimates for the obstacle problem).
Assume that $u^{*},\phi\in H^{1+s}(\Omega)$, $\lambda^{*}=E^{\prime}(u^{*})=-\Delta u^{*}+f\in H^{1-r}(\Omega)$ for fixed $s,r\in(0,1]$ and that $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega)}\lesssim 1$. Moreover, for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair, assume that the mesh is quasi-uniform and satisfies Assumption 4.
For both the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ and $(\mathbb{P}_{1},\mathbb{P}_{1})$ element pairs, the following error estimate holds:
(119) $\|u^{*}-u_{h}^{\ell}\|_{H^{1}(\Omega)}^{2}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{H^{-1}(\Omega)}\lesssim\frac{C_{\mathrm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}\,h^{2\cdot\min\{r,s\}},$
where the constants $C_{\mathrm{stab}}$ and $C_{\mathrm{ref}}$ are independent of $h,\;\ell$, and the proximity parameters $\alpha_{k}$.
In addition, for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, we have the following estimate on the bound-preserving approximation $\tilde{u}_{h}^{\ell}$:
(120) $\|u^{*}-\tilde{u}_{h}^{\ell}\|^{2}_{L^{2}(\Omega)}\lesssim\frac{C_{\mathrm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}\,h^{2\cdot\min\{r,s\}}.$
Proof.
We verify the assumptions of Theorem 3.6. For the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, Lemma 4.1 shows that 76a-76b and 78a in Assumption 3 hold. For the $(\mathbb{P}_{1},\mathbb{P}_{1})$ elements, the Fortin operator is simply the $L^{2}(\Omega)$ projection onto $V_{h}$. It is standard to show that 76a-76b and 78a hold if the mesh is quasi-uniform, as assumed in this case; see, e.g., [24, Proposition 22.21] and [36, Remark 5.2].
To verify Assumption 2, define the enriching map as follows:
(121) $\mathcal{E}_{h}w:=\mathcal{C}_{h}(w-\phi)+\phi$
where $\mathcal{C}_{h}$ is the appropriate quasi-interpolant analyzed in Lemmas 4.2 and 4.3. Observe that $\mathcal{E}_{h}w|_{\partial\Omega}=0$ for any $w\in H^{1}_{0}(\Omega)$, which follows from $\mathcal{C}_{h}w|_{\partial\Omega}=0$ and $(\phi-\mathcal{C}_{h}\phi)|_{\partial\Omega}=0$ since $\phi$ is a constant on $\partial\Omega$. Hence, we obtain that $\mathcal{E}_{h}w\in H^{1}_{0}(\Omega)$ because $\mathcal{C}_{h}(w-\phi)\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}(\Omega)$ and $\phi\in H^{1}(\Omega)$. By construction of the interpolants, cf. 106 and 114, we have that $\mathcal{C}_{h}(w-\phi)(z)\geq 0$ for any $z\in\mathcal{N}_{h}$. In addition, for $z\in\partial\Omega$, $\mathcal{C}_{h}(w-\phi)(z)=-\phi|_{\partial\Omega}\geq 0$. Hence, we conclude that $\mathcal{E}_{h}w\geq\phi$ and $\mathcal{E}_{h}w\in K$.
Observing that $\mathcal{E}_{h}w=\mathcal{C}_{h}w+\varepsilon$ where $\varepsilon=\phi-\mathcal{C}_{h}\phi$, the inequality 78b is verified in Lemma 4.2 and Lemma 4.3. It remains to verify 78c. This follows directly from 83
(122) $\|\varepsilon\|_{H^{t}(\Omega)}$ $=\|\phi-\mathcal{C}_{h}\phi\|_{H^{t}(\Omega)}\lesssim h^{1+s-t}|\phi|_{H^{1+s-t}(\Omega)}.$
Having verified both Assumptions 2 and 3, we invoke Theorem 3.6 to conclude that 119 holds.
To show 120, we utilize Lemma 3.5 and note that $P_{W_{h}}\tilde{u}_{h}^{\ell}=\tilde{u}_{h}^{\ell}$ since $\tilde{u}_{h}^{\ell}\in\mathbb{P}_{0}(\mathcal{T}_{h})$. Therefore, since $B$ is the identity map and $\Omega_{d}=\Omega$ in this case, 71 reads
(123) $\|u^{*}-\tilde{u}_{h}^{\ell}\|_{L^{2}(\Omega)}$ $\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})u_{h}^{\ell}\|_{L^{2}(\Omega)}$
$\leq\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})(u_{h}^{\ell}-u^{*})\|_{L^{2}(\Omega)}+\|(I-P_{W_{h}})u^{*}\|_{L^{2}(\Omega)}$
$\lesssim\|u_{h}^{\ell}-u^{*}\|_{L^{2}(\Omega)}+h|u^{*}|_{H^{1}(\Omega)},$
where we used the stability of $P_{W_{h}}$ and that $\|u^{*}-P_{W_{h}}u^{*}\|_{L^{2}(T)}\lesssim h_{K}|u^{*}|_{H^{1}(T)}$ for any $T\in\mathcal{T}_{h}$. The final estimate 120 is obtained by combining 119 and the final inequality in 123. ∎
5. Application II: The Signorini problem
We consider the following version of the Signorini problem on a two-dimensional domain for simplicity; cf. Example 1.2. Find $u^{*}\in V=H^{1}_{\mathrm{D}}(\Omega)^{2}$ where $H^{1}_{\mathrm{D}}(\Omega):=\{v\in H^{1}(\Omega;\mathbb{R}^{2})\mid v|_{\Gamma_{\rm D}}=0\}$ minimizing the strain energy function
(124) $\min_{u\in K}E(u),\;\;E(u):=\frac{1}{2}\int_{\Omega}\mathsf{C}\,\epsilon(u):\epsilon(u)\,\operatorname{d}\!x-\int_{\Omega}f\cdot u\,\operatorname{d}\!x,$
where $\partial\Omega=\overline{\Gamma_{\mathrm{D}}\cup\Gamma_{\mathrm{T}}}$ and
(125) $K=\{u\in V\mid u\cdot n\leq g\text{ on }\Gamma_{\mathrm{T}}\}.$
Here, the boundary of $\Omega$ consists of two (relatively) open, disjoint subsets $\partial\Omega=\overline{\Gamma_{\rm T}\cup\Gamma_{\rm D}}$ with $|\Gamma_{\rm T}|>0$ and $|\Gamma_{\rm D}|>0$. For further simplicity, we assume that the contact boundary $\Gamma_{\rm T}$ is an open straight line segment and fix a gap function $g\in H^{1}(\Omega)$ with $g|_{\Gamma_{\rm D}}=\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\geq}0$. Recall that in this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n$.
We work with a quasi-uniform mesh $\mathcal{T}_{h}$, and consider the continuous Lagrange spaces on $\Omega$ and on $\Omega_{d}=\Gamma_{\mathrm{T}}$:
(126a) $V_{h}$ $=(\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{\mathrm{D}}(\Omega))^{2},$
(126b) $W_{h}$ $=\{w_{h}\in C(\overline{\Gamma_{\mathrm{T}}})\mid w_{h}|_{E}\in\mathbb{P}_{1}(E)~\text{ for all }\text{edges }E\subset\Gamma_{\mathrm{T}},\;\;w_{h}=0\text{ on }\partial\Gamma_{\mathrm{T}}\}.$
In the above, $\partial\Gamma_{\mathrm{T}}$ denotes the boundary of $\Gamma_{\mathrm{T}}$; i.e., the nodes shared between $\overline{\Gamma_{\mathrm{D}}}$ and $\overline{\Gamma_{\mathrm{T}}}$. To ensure compatibility, note that the degrees of freedom of $\psi_{h}^{k}$ in 10 are set to $\nabla\mathcal{R}(0)$ on the nodes belonging to $\overline{\Gamma_{\mathrm{D}}}\cap\overline{\Gamma_{\mathrm{T}}}$; cf. Remark 3. In what follows, we will denote the normal and tangential components of a vector field $v\in H^{1}(\Omega)^{2}$ by $v_{n}$ and $v_{\tau}$, respectively.
For the considered problem, one can utilize the following choice of Legendre function $\mathcal{R}$:
(127) $\mathcal{R}(u)=(g-u)\ln(g-u)-(g-u),$
which admits the convex conjugate
(128) $\mathcal{R}^{*}(\psi)=\exp(-\psi)+g\psi.$
We provide Algorithm 3, the application of the PG method (Algorithm 1) to this problem for completeness; see also [19, Example 2].
Algorithm 3 The Proximal Galerkin Method for the Signorini Problem
1:input: Initial latent solution guess $\psi_{h}^{0}\in W_{h}$, a sequence of positive proximity parameters $\{\alpha_{k}\}$.
2:Initialize $k=1$.
3:repeat
4: Find $u_{h}^{k}\in V_{h}$ and $\psi_{h}^{k}\in\nabla\mathcal{R}(0)+W_{h}$ such that
(129a) $\alpha_{k}\,(\mathsf{C}\,\epsilon(u^{k}_{h}),\epsilon(v_{h}))+(v_{h}\cdot n,\psi^{k}_{h}-\psi_{h}^{k-1})_{\Gamma_{\mathrm{T}}}$ $=\alpha_{k}\,(f,v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(129b) $(u^{k}_{h}\cdot n,w_{h})_{\Gamma_{\mathrm{T}}}+(\exp(-\psi^{k}_{h}),w_{h})_{\Gamma_{\mathrm{T}}}$ $=(g,w_{h})_{\Gamma_{\mathrm{T}}}$  $~\text{ for all }w_{h}\in W_{h}.$
5: Assign $k\leftarrow k+1$.
6:until a convergence test is satisfied.
Main goal: We derive error estimates for Algorithm 3 in Corollary 3. We also apply the framework presented in Section 3, utilizing Theorem 3.6. To this end, we proceed by constructing Fortin and enriching operators satisfying Assumptions 2 and 3.
Lemma 5.1 (Fortin operator).
Let $\mathcal{T}_{h}$ be quasi-uniform. There exists a map $\tilde{\Pi}_{h}:H^{1}_{\mathrm{D}}(\Omega)\rightarrow V_{h}$ such that
$\int_{\Gamma_{\mathrm{T}}}\tilde{\Pi}_{h}vw_{h}\operatorname{d}\!s=\int_{\Gamma_{\mathrm{T}}}vw_{h}\operatorname{d}\!s~\text{ for all }w_{h}\in W_{h}.$
In addition, the Fortin operator $\Pi_{h}:H^{1}_{\mathrm{D}}(\Omega)^{2}\rightarrow V_{h}$ given by $\Pi_{h}w=(\tilde{\Pi}_{h}w_{n})n+(\mathcal{SZ}_{h}w_{\tau})\tau$ satisfies 42 and the stability and approximation bounds 76a, 76b and 78a stated in Assumption 3.
Proof.
We define $\tilde{\Pi}_{h}$ and show its stability and approximation properties. The stated bounds for the operator $\Pi_{h}$ can then be deduced form the properties of $\tilde{\Pi}_{h}$ and of the Scott-Zhang interpolant. Let $\pi_{h}$ denote the $L^{2}(\Gamma_{\mathrm{T}})$ projection onto $W_{h}$: for $v\in L^{2}(\Gamma_{\mathrm{T}})$, $\pi_{h}v\in W_{h}$ solves
(130) $\int_{\Gamma_{\mathrm{T}}}(\pi_{h}v-v)w_{h}=0~\text{ for all }w_{h}\in W_{h}.$
For the nodes $z$ on $\Gamma_{\mathrm{T}}$ (i.e., $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$), set $\tilde{\Pi}_{h}v(z)=\pi_{h}v(z)$. On the remaining nodes in $\mathcal{N}_{h}$, set $\tilde{\Pi}_{h}v(z)=\mathcal{SZ}_{h}v(z)$. We split $\mathcal{T}_{h}$ into two subsets $\mathcal{T}_{\mathrm{T},h}$ and $\mathcal{T}_{h}\backslash\mathcal{T}_{\mathrm{T},h}$, where $\mathcal{T}_{\mathrm{T},h}$ consists of elements that share a node $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$. Denoting the associated macro-element by $\Delta_{T}$, we observe that
$\|\tilde{\Pi}_{h}v\|_{L^{2}(T)}=\|\mathcal{SZ}_{h}v\|_{L^{2}(T)}\lesssim\|v\|_{L^{2}(\Delta_{T})}+h_{T}\|\nabla v\|_{L^{2}(\Delta_{T})},\quad T\in\mathcal{T}_{h}\backslash\mathcal{T}_{\mathrm{T},h},$
where we used that $\tilde{\Pi}_{h}v(z)=\mathcal{SZ}_{h}v(z)$ for all $z\in\mathcal{N}_{h}\backslash\Gamma_{\mathrm{T}}$ and the local stability properties of $\mathcal{SZ}_{h}$ [46]. For $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$, we can apply a local inverse inequality to deduce that
(131) $|\tilde{\Pi}_{h}v(z)|=|\pi_{h}v(z)|\lesssim h_{E}^{-1/2}\|\pi_{h}v\|_{L^{2}(E)},\quad E\subset\Gamma_{\mathrm{T}}.$
Therefore, since $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}$, we obtain
$\|\tilde{\Pi}_{h}v\|_{L^{2}(T)}^{2}\lesssim\sum_{z\in\mathcal{V}_{T}}\|\tilde{\Pi}_{h}v(z)\varphi_{z}\|_{L^{2}(T)}^{2}\lesssim h_{E}\|\pi_{h}v\|^{2}_{L^{2}(E)}+\|v\|_{L^{2}(\Delta_{T})}^{2}+h_{T}^{2}\|\nabla v\|_{L^{2}(\Delta_{T})}^{2},$
for all $T\in\mathcal{T}_{\mathrm{T},h}$ and some facet $E\subset\Gamma_{\mathrm{T}}$. In particular, if $T$ shares a facet $E$ with $\Gamma_{\rm T}$, then we can select this facet. Otherwise, the facet of a neighboring element is selected. Summing over all the elements and noting that each facet $E\subset\Gamma_{\mathrm{T}}$ is counted at most $\max_{z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}}\operatorname{card}(\omega_{z})$ times, which is bounded uniformly w.r.t. $h$ thanks to shape-regularity of the mesh $\mathcal{T}_{h}$, we obtain that
(132) $\|\tilde{\Pi}_{h}v\|_{L^{2}(\Omega)}\lesssim h^{1/2}\|\pi_{h}v\|_{L^{2}(\Gamma_{\mathrm{T}})}+\|v\|_{L^{2}(\Omega)}+h\|\nabla v\|_{L^{2}(\Omega)}.$
Using the stability of the projection $\pi_{h}$ and a global trace inequality coming from the quasi-uniformity of $\mathcal{T}_{h}$, we obtain that
(133) $h^{1/2}\|\pi_{h}v\|_{L^{2}(\Gamma_{\mathrm{T}})}\leq h^{1/2}\|v\|_{L^{2}(\Gamma_{\mathrm{T}})}\lesssim\|v\|_{L^{2}(\Omega)}+h\|\nabla v\|_{L^{2}(\Omega)}.$
Combining 133 and 132 shows that 76a holds.
We proceed to verify the error estimate 78a. Note that $\Pi_{h}v_{h}=v_{h}$ for any $v_{h}\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{\rm D}(\Omega)$. To see this, note that $\pi_{h}v_{h}=v_{h}|_{\Gamma_{\rm{T}}}$ since $v_{h}|_{\Gamma_{\mathrm{T}}}\in W_{h}$ and $\mathcal{SZ}_{h}v_{h}(z)=v_{h}(z)$ for all $z\in\mathcal{N}_{h}$ [46]. Hence,
(134) $\|v-\tilde{\Pi}_{h}v\|_{L^{2}(\Omega)}$ $\leq\|v-\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}+\|\tilde{\Pi}_{h}(v-\mathcal{SZ}_{h}v)\|_{L^{2}(\Omega)}$
$\lesssim\|v-\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}+h\|\nabla(v-\mathcal{SZ}_{h}v)\|_{L^{2}(\Omega)},$
where we used the just proven stability property 76a of $\Pi_{h}$. The estimate on the $L^{2}$ error stated in 78a can now be concluded from the approximation properties of $\mathcal{SZ}_{h}$. To show 76b, one uses an inverse estimate followed by approximation properties. Namely, we have
(135) $\|\nabla\tilde{\Pi}_{h}v\|_{L^{2}(\Omega)}\leq\|\nabla(\tilde{\Pi}_{h}v-\mathcal{SZ}_{h}v)\|_{L^{2}(\Omega)}+\|\nabla\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}\\ \lesssim h^{-1}\|\tilde{\Pi}_{h}v-\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}+\|\nabla\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}\lesssim\|\nabla v\|_{L^{2}(\Omega)},$
where we used the stability of $\mathcal{SZ}_{h}$ in the $H^{1}(\Omega)$-seminorm, see [8, Corollary 4.8.15]. The final error estimate for $\tilde{\Pi}_{h}$ in the $H^{1}(\Omega)$-seminorm is proven similarly. ∎
In order to check Assumption 2, we observe that the set $K_{h}$ 74 satisfies
(136) $K_{h}\subset\{u_{h}\in V_{h}\mid(u_{h}\cdot n-g,\varphi_{z})_{\Gamma_{\mathrm{T}}}\leq 0~\text{ for all }z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}\}.$
To define the enriching map required by Assumption 2, we begin by defining a modified Clément interpolant ${\mathcal{C}}_{h}:H^{1}(\Omega)\rightarrow H^{1}(\Omega)\cap\mathbb{P}_{1}(\mathcal{T}_{h})$ using the strategy outlined in Section 4.2. To this end, we require a technical assumption on the mesh. Note that this assumption can be easily satisfied by a local refinement near the boundary $\partial\Gamma_{\mathrm{T}}=\overline{\Gamma_{\rm T}}\cap\overline{\Gamma_{\rm D}}$.
Assumption 5.
Assume that $\Gamma_{\mathrm{T}}$ contains at least two mesh facets. For each boundary node $z\in\partial\Gamma_{\mathrm{T}}$, let $E_{z}^{1}$ denote the unique boundary facet in $\Gamma_{\rm T}$ containing $z$. Likewise, let $E_{z}^{2}\subset\Gamma_{T}$ denote the unique boundary facet neighboring $E_{z}^{1}$ and not containing $z$. Assume that $|E_{z}^{1}|\geq|E_{z}^{2}|$.
For any boundary node $z\in\mathcal{N}_{h}\cap\partial\Omega$, we let $\tilde{\omega}_{z}$ denote the union of the boundary facets sharing $z$. Then, denoting the three distinct vertices of the facets in $\tilde{\omega}_{z}$ by $\{z_{0},z_{1},z_{2}\}$ with $z_{0}=z$, we define
(137) ${\mathcal{C}}_{h}v=\sum_{z\in\mathcal{N}_{h}}v_{z}\varphi_{z},\;\;\;v_{z}=\sum_{i=0}^{2}\frac{\alpha_{z,i}}{\int_{\tilde{\omega}_{z_{i}}}\varphi_{z_{i}}}\int_{\tilde{\omega}_{z_{i}}}v\varphi_{z_{i}}\operatorname{d}\!s~\text{ if }z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}},$
where the weights $\alpha_{z,i}\geq 0$, $i=0,1,2$, are defined below. For the remaining nodes in $\mathcal{N}_{h}$, we set ${v}_{z}=\mathcal{SZ}_{h}v(z)$.
We now define ${\alpha}_{z,i}$. Let $(\hat{w}_{j},\hat{x}_{j})$, $j=0,1$, be the non-negative weights and points given by the Gauss–Radau quadrature rule [24, Table 6.1], which is an exact quadrature rule for polynomials of degree $2$ on the reference facet $\hat{E}$. Let $F_{E}$ denote the affine linear transformation from the reference facet $\hat{E}$ to a facet $E\subset\Gamma_{\mathrm{T}}$. Observe that these quadrature points define weights and points $(w_{j},x_{j})$ with $x_{j}\in\tilde{\omega}_{z}$ for $j=0,1,2,3$, such that for any continuous piecewise second-order polynomial $v$ on $\tilde{\omega}_{z}=E_{1}\cup E_{2}$, we can write
(138) $\int_{\tilde{\omega}_{z}}v\operatorname{d}\!s=\sum_{j=0}^{1}\frac{|E_{1}|}{|\hat{E}|}\hat{w}_{j}v(F_{E_{1}}(\hat{x}_{j}))+\sum_{j=0}^{1}\frac{|E_{2}|}{|\hat{E}|}\hat{w}_{j}v(F_{E_{2}}(\hat{x}_{j}))=\sum_{j=0}^{3}w_{j}v(x_{j}).$
For any $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$, there exists a point $s_{z}\in\tilde{\omega}_{z}$ such that
(139) $s_{z}=\frac{1}{\int_{\tilde{\omega}_{z}}\varphi_{z}}\int_{\tilde{\omega}_{z}}x\varphi_{z}(x)\operatorname{d}\!s=\frac{1}{\int_{\tilde{\omega}_{z}}\varphi_{z}}\sum_{j=0}^{3}w_{j}x_{j}\varphi_{z}(x_{j}).$
Therefore, $s_{z}$ is a convex combination of the points $\{x_{j}\}_{j=0,\ldots,3}$ belonging to $\tilde{\omega}_{z}$. Note that $z$ must lie either on the line connecting $s_{z_{0}}$ to $s_{z_{1}}$ or on the line connecting $s_{z_{0}}$ to $s_{z_{2}}$. This means that there exist convex weights $\alpha_{z_{i}}\geq 0$ (with one $\alpha_{z_{i}}=0$), such that
(140) $z=\sum_{i=0}^{2}\alpha_{z_{i}}s_{z_{i}},\quad\sum_{i=0}^{2}\alpha_{z_{i}}=1,\quad\alpha_{z_{i}}\geq 0.$
Further, Assumption 5 guarantees that the weights corresponding to the boundary nodes in $\partial\Gamma_{\mathrm{T}}$ are zero in 140; i.e., $\alpha_{z}=0$ if $z\in\partial\Gamma_{\mathrm{T}}$. In turn, for each $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$, there exists an element $T_{z}\in\mathcal{T}_{h}$, $T_{z}\subset\cup_{z_{i}}\omega_{z_{i}}$, such that if $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$, it holds that
(141) ${\mathcal{C}}_{h}p(z)=p(z).$
Lemma 5.2 (Approximation properties of $\mathcal{C}_{h}$).
Let $\mathcal{T}_{h}$ be quasi-uniform and let Assumption 5 hold. For $0\leq s\leq 1$, it follows that
(142) $\|w-\mathcal{C}_{h}w\|_{L^{2}(\Omega)}+h\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(\Omega)}$ $\lesssim h^{1+s}|w|_{H^{1+s}(\Omega)}$  $~\text{ for all }w\in H^{1+s}(\Omega).$
Proof.
We first consider the error $(\mathcal{C}_{h}v_{h}-v_{h})$ for any $v_{h}\in V_{h}$. For a node $z\in\Gamma_{\mathrm{T}}$, we consider the element $T_{z}$ such $\mathcal{C}_{h}p(z)=p(z)$ for $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$; cf. 141. Then,
$(\mathcal{C}_{h}v_{h}-v_{h})(z)=\mathcal{C}_{h}(v_{h}-p)(z)+(p-v_{h})(z),\;\;p\in\mathbb{P}_{1}(\Delta_{T_{z}}).$
With Cauchy–Schwarz inequality, the observation that $\|\varphi_{z}\|_{L^{2}(\omega_{z_{i}})}\lesssim h^{1/2}$, and a trace inequality, we estimate
(143) $|\mathcal{C}_{h}(v_{h}-p)(z)|$ $\leq\sum_{i=0}^{2}|\tilde{\alpha}_{z_{i}}|(2|\tilde{\omega}_{z_{i}}|^{-1})\|v_{h}-p\|_{L^{2}(\tilde{\omega}_{z,i})}\|\varphi_{z}\|_{L^{2}(\tilde{\omega}_{z_{i}})}$
$\lesssim\sum_{i=0}^{2}|\tilde{\alpha}_{z_{i}}|h^{-1/2}\|v_{h}-p\|_{L^{2}(\tilde{\omega}_{z_{i}})}$
$\lesssim h^{-1}\|v_{h}-p\|_{L^{2}(\Delta_{T_{z}})}.$
In addition, we apply an inverse estimate to bound
(144) $|(p-v_{h})(z)|\leq\|p-v_{h}\|_{L^{\infty}(\Delta_{T_{z}})}\lesssim h^{-1}\|p-v_{h}\|_{L^{2}(\Delta_{T_{z}})}$
since $(p-v_{h})\in V_{h}$. Combining the above estimates, followed by applying the triangle inequality and the Bramble–Hilbert Lemma, we arrive at
(145) $|(\mathcal{C}_{h}v_{h}-v_{h})(z)|$ $\lesssim h^{-1}(\|v_{h}-v\|_{L^{2}(\Delta_{T_{z}})}+\|v-p\|_{L^{2}(\Delta_{T_{z}})})$
$\lesssim h^{-1}\|v_{h}-v\|_{L^{2}(\Delta_{T_{z}})}+h^{s}|v|_{H^{1+s}(\Delta_{T_{z}})}.$
Select $v_{h}=\mathcal{SZ}_{h}v$ and recall that by definition of $\mathcal{C}_{h}$, we have that $(\mathcal{C}_{h}-\mathcal{SZ}_{h})v(z)=0$ for all $z\in\mathcal{N}_{h}\backslash\Gamma_{\mathrm{T}}$. Therefore, for any $T\in\mathcal{T}_{h}$, we conclude that
(146) $(\mathcal{C}_{h}v-\mathcal{SZ}_{h}v)|_{T}=\sum_{z\in\Gamma_{\mathrm{T}}\cap\mathcal{V}_{T}}(\mathcal{C}_{h}v-\mathcal{SZ}_{h}v)(z)\varphi_{z}.$
Using that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}$, 145, and applying Cauchy–Schwarz inequality, we obtain that
(147) $\|\mathcal{C}_{h}v-\mathcal{SZ}_{h}v\|_{L^{2}(T)}\lesssim\sum_{z\in\Gamma_{\mathrm{T}}\cap\mathcal{V}_{T}}\left(h^{1+s}|v|_{H^{1+s}(\Delta_{T_{z}})}+\|\mathcal{SZ}_{h}v-v\|_{L^{2}(\Delta_{T_{z}})}\right).$
Summing over all the elements that contain a node in $\Gamma_{\mathrm{T}}$ and applying an inverse estimate, we find
(148) $\|\mathcal{C}_{h}v-\mathcal{SZ}_{h}v\|_{L^{2}(\Omega)}+h\|\nabla(\mathcal{C}_{h}v-\mathcal{SZ}_{h}v)\|_{L^{2}(\Omega)}\lesssim h^{1+s}|v|_{H^{1+s}(\Omega)}+\|\mathcal{SZ}_{h}v-v\|_{L^{2}(\Omega)}.$
The result can be concluded by applying the triangle inequality and the approximation properties of $\mathcal{SZ}_{h}$ [46]. ∎
Corollary 3 (A priori error estimate for the Signorini problem).
Let $u^{*}$ solve 124 and let $(u_{h}^{\ell},\psi_{h}^{\ell})$ come from Algorithm 3. Assume that $u^{*}\in H^{1+s}(\Omega)^{2}$, $g\in H^{1+s}(\Omega)$ and $\mathrm{div}(\mathsf{C}\epsilon(u))-f\in H^{-1+r}(\Omega)^{2}$ for some $r,s\in(0,1]$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Gamma_{\rm T})}\lesssim 1$, and Assumption 5 holds. The following error estimate holds:
(149) $\|u^{*}-u_{h}^{\ell}\|_{H^{1}(\Omega)}^{2}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{H^{-1/2}(\Gamma_{\mathrm{T}})}\leq\frac{C_{\mathrm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}\,h^{2\cdot\min\{r,s\}}.$
Proof.
Given the Fortin operator that we constructed in Lemma 5.1, we are in the setting of Section 3. We only need to verify the assumptions of Theorem 3.6. To this end, we define the normal and tangential components of $\mathcal{E}_{h}$ as follows:
(150) $(\mathcal{E}_{h}w)_{n}=\mathcal{C}_{h}(w_{n}-g)+g,\quad(\mathcal{E}_{h}w)_{\tau}=\mathcal{SZ}_{h}w_{\tau},$
To check Assumption 2, we first note that $(\mathcal{E}_{h}w)_{n}=0$ on $\Gamma_{\rm D}$, $(\mathcal{E}_{h}w)_{n}\in H_{\mathrm{D}}^{1}(\Omega)$ and $(\mathcal{E}_{h}w)_{\tau}\in H_{\mathrm{D}}^{1}(\Omega)$ from the properties of $\mathcal{SZ}_{h}$. This implies that $\mathcal{E}_{h}w\in H^{1}_{\mathrm{D}}(\Omega)^{2}$. We now check that $g-\mathcal{E}_{h}w\cdot n\geq 0$ on $\Gamma_{\mathrm{T}}$ to ensure that $\mathcal{E}_{h}w\in K$. For $w\in K\cup K_{h}$, we have that $(u_{h}\cdot n-g,\varphi_{z})_{\tilde{\omega}_{z}}\leq 0$ for any $z\in\Gamma_{\mathrm{T}}$ since the support of $\varphi_{z}$ on $\Gamma_{\mathrm{D}}$ is $\tilde{\omega}_{z}$. The additional Assumption 5 on the mesh guarantees that $\tilde{\omega}_{z}$ for $z\in\partial\Gamma_{\rm T}$ is not included in the definition of $\mathcal{C}_{h}$ 137. From 137, we now obtain that $\mathcal{C}_{h}(w_{n}-g)(z)\leq 0$ for all $z\in\Gamma_{\mathrm{T}}$. In addition, for the nodes on $\partial\Gamma_{\rm T}$, we find $\mathcal{C}_{h}(w_{n}-g)=-g|_{\Gamma_{\rm D}}=-\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}0$. Since $g-\mathcal{E}_{h}w\cdot n=-\mathcal{C}_{h}(w_{n}-g)$, we conclude that $(g-\mathcal{E}_{h}w\cdot n)\geq 0$ on $\Gamma_{\mathrm{T}}$. This verifies Assumption 2.
Noting that $\mathcal{E}_{h}w=(\mathcal{C}_{h}w_{n},\mathcal{SZ}_{h}w_{\tau})+(g-\mathcal{C}_{h}g,0)$, Assumption 3 is verified by applying Lemma 5.1, Lemma 5.2, standard estimates on $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon=(g-\mathcal{C}_{h}g,0)$} similar to 122, and the approximation properties of the Scott–Zhang interpolant [46]. Details are skipped for brevity. ∎
6. Conclusion
The PG method offers a versatile and efficient approach for solving variational problems with pointwise inequality constraints. We provided an abstract framework for its a priori error analysis in the context of quadratic optimization problems with such constraints. We utilized this framework to derive optimal error estimates for the obstacle and Signorini problems, demonstrating the effectiveness of the PG method. Numerically, the PG method achieves high-order error rates; however, the error analysis we presented is currently limited to the lowest-order conforming approximations. Future work could extend our results to high-order approximation spaces and a broader class of energy functionals.
Acknowledgments
We thank Ioannis P. A. Papadopoulos for his comments on the manuscript and Thomas M. Surowiec and Dohyun Kim for numerous interesting discussions. BK and RM were supported in part by the U.S. Department of Energy Office of Science, Early Career Research Program under Award Number DE-SC0024335 and by the Center for Information Geometric Mechanics (CIGMO), a PSAAP-IV Focused Investigatory Center funded by the Department of Energy under Award Number DE-NA0004261. Keith’s work was also supported in part by the Alfred P. Sloan Foundation via a Sloan Research Fellowship in Mathematics. MZ acknowledges support from an ETH Postdoctoral Fellowship for the project “Reliable, Efficient, and Scalable Methods for Scientific Machine Learning”.
Appendix A A stability result
In this section, we show a stability result (Theorem A.1 below) for the discrete iterates $(u_{h}^{k},\psi_{h}^{k})$. For simplicity, we assume that $\nabla\mathcal{R}^{*}$ takes the following form:
(151a) $\nabla\mathcal{R}^{*}(\psi)(x)=\phi_{0}(x)\nabla\mathcal{R}^{*}_{0}(\psi(x))+\phi_{1}(x),$
where $\phi_{0}\in L^{\infty}(\Omega_{d})$ with $\phi_{0}(x)\geq\underline{\phi_{0}}>0$ a.e. in $\Omega$, $\phi_{1}\in Q$ and
(151b) $\mathcal{R}_{0}^{*}(\psi)-\nu_{1}|\psi|\geq c_{1}~\text{ for all }\psi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m}),$
for some $\nu_{1}\geq 0$ and $c_{1}\in\mathbb{R}$.
The generalized Shannon entropy 30 does not fit into the form 151 for all obstacles $\phi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. However, the decomposition 151 is written with sufficient generality to accommodate the settings analyzed in this paper. In particular, consider the obstacle problem from Section 4, where $\phi\in H^{1}(\Omega)$ and $\phi|_{\partial\Omega}=-\delta$ for a positive constant $\delta$. In this case, 151 holds with $\mathcal{R}_{0}^{*}(\psi)=\exp(\psi)-\delta\psi$, $\phi_{0}=1$, and $\phi_{1}=\phi+\delta$, implying $\nu_{1}=\delta$. For the Fermi–Dirac entropy given in Example 2.2, which is a suitable choice for bilateral obstacle problems ($B=\mathrm{id}$ and $Q=V=H^{1}_{0}(\Omega)$), we can write
(152) $\nabla\mathcal{R}^{*}(\psi)=\frac{1}{2}(\overline{u}-\underline{u})\frac{\exp(\psi)-1}{\exp(\psi)+1}+\frac{1}{2}(\overline{u}+\underline{u}).$
Thus, defining $\mathcal{R}_{0}^{*}(\psi)=2\ln(\exp(\psi)+1)-\psi$ and assuming that $\overline{u}-\underline{u}\geq\underline{\phi_{0}}$ and $\overline{u}+\underline{u}\in H^{1}_{0}(\Omega)$, we deduce that 151 holds with $\nu_{1}=1$. Notably, if $\underline{u},\overline{u}\in\mathbb{R}$ with $\underline{u}<0<\overline{u}$, the condition $\underline{u}+\overline{u}\in H^{1}_{0}(\Omega)$ is not generally satisfied. However, in this case, we can still verify 151 using an alternative decomposition given by $\mathcal{R}_{0}^{*}(\psi)=(\overline{u}-\underline{u})\ln(\exp(\psi)+1)+\underline{u}\psi$, $\phi_{0}=1$, and $\phi_{1}=0$. The Hellinger entropy, introduced in Example 2.3, readily satisfies 151 with $\mathcal{R}_{0}^{*}(\psi)=\gamma\sqrt{1+|\psi|^{2}}$, $\phi_{0}=1$, and $\phi_{1}=0$, implying $\nu_{1}=\gamma$ and $c_{1}=0$.
Theorem A.1 (Stability).
Assume that 151 holds. Further, assume that $\psi_{h}^{0}$ is selected so that $\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}}\lesssim 1$. Then there exists a constant $C_{\mathrm{stab}}$, independent of $h$, $\ell$, and $\{\alpha_{k}\}_{k=2,\ldots}$, such that
(153) $\nu\|u_{h}^{\ell}\|_{V}+\|\lambda_{h}^{\ell}\|_{Q^{\prime}}+\frac{\|\psi_{h}^{\ell}\|_{Q^{\prime}}+\nu_{1}\underline{\phi_{0}}\|\psi_{h}^{\ell}\|_{L^{1}(\Omega)}}{\sum_{k=1}^{\ell}\alpha_{k}}\leq C_{\mathrm{stab}}~\text{ for all }\ell\geq 1\text{ and }h>0.$
Proof.
We first observe that testing 10a with $v_{h}=u_{h}^{k}$ and 10b with $w_{h}=\psi_{h}^{k}$, subtracting the resulting equations, and using the expression 151a yields
(154) $\alpha_{k}a(u_{h}^{k},u_{h}^{k})+(\phi_{0}\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega_{d}}=\alpha_{k}F(u_{h}^{k})+b(u_{h}^{k},\psi_{h}^{k-1})-(\phi_{1},\psi_{h}^{k})_{\Omega_{d}}$
for every $k\geq 1$. To handle the second term above, we split $\Omega_{d}$ into $\Omega_{d}^{-}=\{x\in\Omega_{d}\mid\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k})(x)\psi_{h}^{k}(x)\leq 0\}$ and $\Omega_{d}^{+}=\{x\in\Omega_{d}\mid\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k})(x)\psi_{h}^{k}(x)\geq 0\}$. Proceeding, we estimate
$(\phi_{0}\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega^{+}_{d}}$ $\geq\underline{\phi_{0}}(\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega^{+}_{d}}$,
$(\phi_{0}\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega^{-}_{d}}$ $\geq\|\phi_{0}\|_{L^{\infty}(\Omega_{d})}(\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega^{-}_{d}}.$
Using the subgradient inequality and (151b), we also bound
(155) $(\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega_{d}^{\pm}}\geq(\mathcal{R}_{0}^{*}(\psi_{h}^{k})-\mathcal{R}_{0}^{*}(0),1)_{\Omega_{d}^{\pm}}\geq\nu_{1}\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d}^{\pm})}+(c_{1}-\mathcal{R}_{0}^{*}(0),1)_{\Omega_{d}^{\pm}}.$
Noting that $c_{1}\leq\mathcal{R}_{0}^{*}(0)$ by 151b, the above estimates yield
(156) $(\phi_{0}\nabla\mathcal{R}_{0}^{*}(\psi_{h}^{k}),\psi_{h}^{k})_{\Omega_{d}}\geq\nu_{1}\underline{\phi_{0}}\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}+\|\phi_{0}\|_{L^{\infty}(\Omega_{d})}(c_{1}-\mathcal{R}_{0}^{*}(0),1)_{\Omega_{d}}.$
Employing the coercivity of the bilinear form $a$ 8, we obtain that
(157) $\alpha_{k}\nu\|u_{h}^{k}\|_{V}^{2}+\nu_{1}\underline{\phi_{0}}\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}\leq(\alpha_{k}\|F\|_{V^{\prime}}+\|B^{\prime}\psi_{h}^{k-1}\|_{V^{\prime}})\|u_{h}^{k}\|_{V}+\|\psi_{h}^{k}\|_{Q^{\prime}}\|\phi_{1}\|_{Q}+c_{2},$
where $c_{2}:=\|\phi_{0}\|_{L^{\infty}(\Omega_{d})}(\mathcal{R}_{0}^{*}(0)-c_{1},1)_{\Omega_{d}}\geq 0$.
Proceeding, we obtain a bound on $\|\psi_{h}^{k}\|_{Q^{\prime}}$. Using the Fortin operator 42 together with 10a, we write
(158) $\langle B^{\prime}\psi_{h}^{k},v\rangle=b(v,\psi_{h}^{k})=b(\Pi_{h}v,\psi_{h}^{k})=b(\Pi_{h}v,\psi_{h}^{k-1})-\alpha_{k}a(u_{h}^{k},\Pi_{h}v)+\alpha_{k}F(\Pi_{h}v)$
for any $v\in V$. The surjectivity of $B:V\rightarrow Q$ and the continuity of the bilinear form $a$ and operator $\Pi_{h}$ give
(159) $\beta\|\psi_{h}^{k}\|_{Q^{\prime}}\leq\|B^{\prime}\psi_{h}^{k}\|_{V^{\prime}}\leq\|\Pi_{h}\|(\|B^{\prime}\psi_{h}^{k-1}\|_{V^{\prime}}+\alpha_{k}M\|u_{h}^{k}\|_{V}+\alpha_{k}\|F\|_{V^{\prime}}),$
where $\beta>0$ is the same constant appearing in the LBB condition 41. Using the above bound in 157 and appropriate applications of Young’s inequality shows that there exists a constant $M_{k}$ (depending only on $\alpha_{k}$, $M$, $\|F\|_{V^{\prime}}$, $\|\Pi_{h}\|$, $\|\phi_{1}\|_{Q}$, $\beta$, and $c_{2}$) such that
(160) $\nu\|u_{h}^{k}\|_{V}^{2}+\nu_{1}\underline{\phi_{0}}\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}\leq M_{k}+\|B^{\prime}\psi_{h}^{k-1}\|^{2}_{V^{\prime}}.$
Selecting $k=1$ in the above estimate, recalling that $\|\Pi_{h}v\|_{V}\lesssim\|v\|_{V}$, and using the assumption that $\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}}\lesssim 1$ for all $v\in V$ implies that there is a constant $C_{1}$ (independent of $h$ and $\alpha_{k}$ for $k>1$) such that
(161) $\nu\|u_{h}^{1}\|^{2}_{V}\leq M_{1}+\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}}^{2}:=C_{1}^{2}.$
To obtain a bound on $\|u_{h}^{\ell}\|_{V}$ for $\ell>1$, we use the energy dissipation property (i.e., Lemma 3.2):
(162) $\frac{1}{2}a(u_{h}^{\ell},u_{h}^{\ell})-F(u_{h}^{\ell})=E(u_{h}^{\ell})\leq E(u_{h}^{1}).$
Using 8, we obtain that
(163) $\frac{\nu}{2}\|u_{h}^{\ell}\|_{V}^{2}$ $\leq M\|u_{h}^{1}\|_{V}^{2}+\|F\|_{V^{\prime}}(\|u_{h}^{1}\|_{V}+\|u_{h}^{\ell}\|_{V})$
$\leq MC_{1}^{2}+\|F\|_{V^{\prime}}C_{1}+\frac{1}{\nu}\|F\|_{V^{\prime}}^{2}+\frac{\nu}{4}\|u_{h}^{\ell}\|_{V}^{2}.$
We can then conclude that
(164) $\frac{\nu}{4}\|u_{h}^{\ell}\|_{V}^{2}\leq MC_{1}^{2}+\|F\|_{V^{\prime}}C_{1}+\frac{1}{\nu}\|F\|_{V^{\prime}}^{2}:=C_{2}^{2}.$
Employing again that $B:V\rightarrow Q$ is surjective, we obtain that
(165) $\beta\|\lambda_{h}^{\ell}\|_{Q^{\prime}}\leq\|B^{\prime}\lambda_{h}^{\ell}\|_{V^{\prime}}\leq M\|u_{h}^{\ell}\|_{V}+\|F\|_{V^{\prime}},$
where we used (10a) and similar arguments to (159). The bounds 164 and 165 yield the first two terms in 153.
To show the bound on $\|\psi_{h}^{\ell}\|_{Q^{\prime}}$, we define the weighted averages $\overline{u}_{h}^{\ell}=\sum_{k=1}^{\ell}\alpha_{k}u_{h}^{k}/\sum_{k=1}^{\ell}\alpha_{k}$ and sum 10a from $k=1$ to $k=\ell$. This gives
(166) $a(\overline{u}_{h}^{\ell},v_{h})+\frac{1}{\sum_{k=1}^{\ell}\alpha_{k}}b(v_{h},\psi_{h}^{\ell})=F(v_{h})+\frac{1}{\sum_{k=1}^{\ell}\alpha_{k}}b(v_{h},\psi_{h}^{0})~\text{ for all }v_{h}\in V_{h}.$
We now observe that for any $v\in V$,
(167) $b(v,\psi_{h}^{\ell})=b(\Pi_{h}v,\psi_{h}^{\ell})=\left(\sum_{k=1}^{\ell}\alpha_{k}\right)(F(\Pi_{h}v)-a(\overline{u}_{h}^{\ell},\Pi_{h}v))+b(\Pi_{h}v,\psi_{h}^{0}).$
With similar arguments as before and the observation that $\|\overline{u}_{h}^{\ell}\|_{V}\leq 2\nu^{-1/2}C_{2}$, this implies that
(168) $\|B^{\prime}\psi_{h}^{\ell}\|_{V^{\prime}}\leq C_{3}\left(\sum_{k=1}^{\ell}\alpha_{k}+1\right)~\text{ for all }\ell\geq 1,$
where $C_{3}=(\|F\|_{V^{\prime}}+2\nu^{-1/2}MC_{2}+\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}})\|\Pi_{h}\|$. To derive the bound on $\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}$, we substitute 168 into 157. ∎